### v1

# MultiNash-PF: A Particle Filtering Approach for Computing Multiple Local Generalized Nash Equilibria in Trajectory Games

Abstract
Modern-world robotics involves complex environments where multiple autonomous agents must interact with each other and other humans. This necessitates advanced interactive multi-agent motion planning techniques. Generalized Nash equilibrium(GNE), a solution concept in constrained game theory, provides a mathematical model to predict the outcome of interactive motion planning, where each agent needs to account for other agents in the environment. However, in practice, multiple local GNEs may exist. Finding a single GNE itself is complex as it requires solving coupled constrained optimal control problems. Furthermore, finding all such local GNEs requires exploring the solution space of GNEs, which is a challenging task. This work proposes the MultiNash-PF framework to efficiently compute multiple local GNEs in constrained trajectory games. Potential games are a class of games for which a local GNE of a trajectory game can be found by solving a single constrained optimal control problem. We propose MultiNash-PF that integrates the potential game approach with implicit particle filtering, a sample-efficient method for non-convex trajectory optimization. We first formulate the underlying game as a constrained potential game and then utilize the implicit particle filtering to identify the coarse estimates of multiple local minimizers of the game’s potential function. MultiNash-PF then refines these estimates with optimization solvers, obtaining different local GNEs. We show through numerical simulations that MultiNash-PF reduces computation time by up to 50% compared to a baseline approach.
I Introduction
Modern-world robotics applications are inherently multi-agent in nature. For instance, autonomous cars on highways need to account for other vehicles while planning their motion. Many robotics applications require motion planning for selfish agents sharing a common operational space, such as autonomous cars and delivery robots. In such settings, autonomous agents should account for other agents in the shared environment while achieving their individual goals. Autonomous agents need to reason about the impact of their actions on other agents. Game theory provides a formal way to deal with such interactive scenarios [1]. Furthermore, recent works have shown that game-theoretic trajectory planning makes robots more socially acceptable to humans [2].
To model the outcome of such interactive scenarios, Nash equilibrium is a key concept in game theory. When multiple agents- such as robots and humans- interact, a Nash equilibrium corresponds to a joint action such that no agent has incentives to deviate from its equilibrium actions unilaterally. This allows us to account for agents’ individual preferences when interacting in a common environment. We formulate a multi-agent interactive trajectory planning problem as a trajectory game. In robotics, seeking Nash equilibria of the underlying trajectory games has been proven to be effective [3, 4]. Furthermore, in real-world applications, agents may have to satisfy some safety constraints, such as avoiding collisions with obstacles and other agents in the environment. In such constrained multi-agent interactive settings, the interaction outcomes are captured through generalized Nash equilibrium(GNE), which are constrained Nash equilibria. Several recent works have shown great success in utilizing GNE for multi-agent motion planning in interactive domains [5, 6].
However, in practice, multiple local GNEs may exist for a trajectory game. For example, in the two-player trajectory game illustrated in Fig. 1, the agents need to move towards their goals while avoiding collisions. Here, two local GNEs exist: clockwise trajectories and counterclockwise trajectories. If the two agents were to act independently, they could reach different GNEs, which would result in a collision. Therefore, agents need a coordination mechanism to avoid such scenarios. However, identifying all the possible local GNEs is required to enable effective coordination. Therefore, computing multiple local GNEs is critical to aligning players’ preferences over different equilibria [7] or helping coordinate players’ actions via recommendation [8]. Furthermore, different local GNEs correspond to different modes of interactions in multi-agent interactive settings. Therefore, it is essential to efficiently recover all the possible multi-modal outcomes of multi-agent interactions.
Computing different local GNEs requires an efficient way to solve games. In practice, computing even a single GNE is often challenging [9] because solving for GNE amounts to solving a set of coupled constrained optimal control problems. Typically, such a solution requires solving first-order KKT conditions of the underlying problems. Therefore, solving the GNE often reduces to addressing a set of complementarity conditions, a process known to be computationally demanding, especially for large-scale applications [10]. Furthermore, computing all the different local GNEs requires initializing the game solvers with many different initial guesses of the equilibrium trajectories from prior knowledge of the specific scenario. Further, as the number of participating agents increases, the dimension of the nonconvex optimization grows to interactable levels. This further increases the required number of initial guesses for the game solvers to find all local GNEs. To our best knowledge, there is no principled approach to compute multiple local Nash equilibria in constrained trajectory games other than exhaustive random initializations of the game solvers.
Figure 1: The Nash equilibrium trajectories found by MultiNash-PF when two unicycle agents change their positions. First, we use the implicit particle filter to discover the coarse estimates of two different equilibria, and then utilize them as a warm starting for IPOPT to obtain the two Nash equilibrium trajectories.
We propose a novel framework named MultiNash-PF to efficiently compute multiple local GNEs in constrained trajectory games. The approach combines
  1. 1.
    constrained potential games [11], a class of games for which we can compute local Nash equilibria by optimizing a single constrained optimal control problem, and,
  2. 2.
    constraint aware implicit particle filtering, which is a sample-efficient particle filtering method recently used for nonconvex trajectory optimization [12].
Using constrained potential games, it is possible to compute a local GNE by solving a constrained optimal control problem using off-the-shelf optimization solvers. These solvers outperform state-of-the-art constrained dynamic game solvers in computation speed during [13]. After formulating the constrained trajectory game as a constrained potential game, our approach employs the constraint-aware implicit particle filtering method to identify multiple local minima of the potential game. This method eliminates the requirements of random initialization of the optimization solver by efficiently obtaining coarse estimates of different local GNEs. Next, we use the identified minima as initialization for the optimization solver. The solver then computes all the different local GNEs of the original game. The framework of MultiNash-PF is depicted in Fig. 1. Leveraging the sample efficiency of the implicit particle filtering method, MultiNash-PF can explore and identify the multiple local GNEs faster than the random initialization of optimization solvers.
Figure 2: MultiNash-PF identifies all the local GNE trajectories for a challenging two agent trajectory game where two agents swapped positions while avoiding an obstacle. In this scenario, the obstacle radius is greater than the collision avoidance radius of the two agents, leading to the discovery of 6 distinct Nash equilibria. MultiNash-PF effectively obtains coarse estimates of local GNEs through implicit particle filtering, which are then refined using IPOPT to obtain local GNEs.
II Related Works
Game Theoretic Planning. One of the earliest works laying the foundation of game theoretic motion planning was [14]. Game theoretic motion planners that seek Stackelberg equilibria were proposed in [15, 3]. Later, methods that seek Nash equilibria were developed. These methods differ mainly in the way they compute Nash equilibria. Iterative best response methods for computing Nash equilibria were developed in [16, 17]. Iterative linear quadratic approximations to compute Nash equilibria were introduced in [18]. ALGames was developed in [19] which is an augmented lagrangian-based solver for constrained dynamic games. A sequential linear-quadratic technique game-based technique to compute GNE was proposed in [20]. However, these methods are computationally expensive for real-time implementations as the number of agents increases. Potential games-based methods were developed later in [11, 21, 13, 22] to reduce the complexity of computing equilibria. However, the shortcomings of all these methods lie in the fact that they find only one Nash equilibrium of the underlying game and ignore the multi-modal nature of the problem that is crucial for coordination among agents. In [7], the authors recognize the existence of multiple equilibria and focus on inferring the equilibrium of other agents in the environment. However, the methodology relies on computing the different approximate local Nash by rollouts of randomly selected initial strategies, which we show in this work to be computationally intractable with the risk of not identifying all possible equilibria.
Multiple Local Solutions to Constrained Optimal Control. In general, optimal control problems are often formulated as nonconvex constrained optimization problems, which are solved through the conventional gradient-based approaches such as inter-point methods [23] and sequential quadratic programming [24]. However, given an initialization, they can only find a single local optimal solution. However, in the present context of identifying all potential local GNEs, the gradient-based method would require numerous initializations to find all the local solutions. Departing from gradient-based methods, a Bayesian inference approach has been proposed to solve model predictive control problem [25, 26, 27]. The approach formulates the optimal control problem from a Bayesian inference perspective and employs particle filtering as its estimation method. At its core, particle filtering utilizes sequential Monte Carlo sampling to perform state estimation for highly nonlinear systems [28, 29]. The efficiency and accuracy of the estimation largely depend on the quality of the sampled particles. Recently, in [30], an efficient particle filtering method was proposed based on the principle of implicit sampling. This sampling approach relies on the premise that a few particles are required for accurate estimation if the particles are sampled from high-probability regions of the probability distribution function. In addition, the implicit sampling technique can also mitigate the need for a resampling step to avoid filter collapse.
III Multi-agent Constrained Trajectory Games
We formulate the multi-agent interactive trajectory planning problem as an $N$-agent constrained trajectory game. Let $\mathcal{N}:=\{1,2,\ldots,N\}$ denote the set of agents’ indices. We assume that the game is played over a finite time horizon of $\tau\in\mathbb{N}$ time-steps, where $\mathbb{N}$ is the set of natural numbers.
Let $x^{i}_{t}\in\mathbb{R}^{n_{i}}$ and $u_{t}^{i}\in\mathbb{R}^{m_{i}}$ denote the state and control of the $i$-th agent at time $t$, respectively. We assume that the state of the $i$-th agent evolves as follows:
$x_{t+1}^{i}=f^{i}(x_{t}^{i},u_{t}^{i}),$ (1)
for all $t\leq\tau$, where $f^{i}:\mathbb{R}^{n_{i}}\times\mathbb{R}^{m_{i}}\to\mathbb{R}^{n_{i}}$ is a continuously differentiable function characterizing the dynamics of the $i$-th agent. Let $n=\sum_{i=1}^{N}n_{i},m=\sum_{i=1}^{N}m_{i}$. We denote the joint state, input, and dynamics of all agents as $x_{t}\coloneqq\begin{bmatrix}(x_{t}^{1})^{\top}&(x_{t}^{2})^{\top}&\cdots&(x_{t}^{N})^{\top}\end{bmatrix}^{\top}$, $u_{t}\coloneqq\begin{bmatrix}(u_{t}^{1})^{\top}&(u_{t}^{2})^{\top}&\cdots&(u_{t}^{N})^{\top}\end{bmatrix}^{\top}$, and $f\coloneqq\begin{bmatrix}(f^{1})^{\top}&(f^{2})^{\top}&\cdots&(f^{N})^{\top}\end{bmatrix}^{\top}$ respectively. Note that each $f^{i}$ is a vector-valued dynamics function for each agent. Therefore, $f:\mathbb{R}^{n}\times\mathbb{R}^{m}\to\mathbb{R}^{n}$ will be a vector-valued dynamics function for the joint system.
In constrained trajectory games, each agent has to satisfy some constraints that couple different agents’ states and controls and, in general, depend on other agents’ states and actions. Let $g:\mathbb{R}^{n}\times\mathbb{R}^{m}\to\mathbb{R}^{c}$ with $c\in\mathbb{N}$ denote a continuously differentiable function that defines the constraints on the joint state $x_{t}$ and joint control $u_{t}$ at each time:
$g(x_{t},u_{t})\leq{0}_{c},$ (2)
where ${0}_{c}$ is a vector of zeros in $\mathbb{R}^{c}$. The aim of the constraint function $g$ is to capture individual agents’ strict preferences, such as constraints for each agent avoiding obstacles, or avoiding collision with other agents.
We assume that each agent $i$ aims to minimize an objective function. Let $Q^{i}$ and $Q_{\tau}^{i}$ be positive semidefinite and $R^{i}$ be positive definite matrices for all $i\in\mathcal{N}$. We assume that each agent $i$ seeks to minimize the following objective:
$\frac{1}{2}\left\lVert x_{\tau}^{i}-\hat{x}_{\tau}^{i}\right\rVert_{Q^{i}_{\tau}}^{2}+\frac{1}{2}\sum_{t=0}^{\tau-1}\left\lVert x_{t}^{i}-\hat{x}_{t}^{i}\right\rVert_{Q^{i}}^{2}+\frac{1}{2}\sum_{t=0}^{\tau}\left\lVert u_{t}^{i}\right\rVert_{R^{i}}^{2}$ (3)
where $\hat{x}_{t}^{i}\in\mathbb{R}^{n_{i}}$ is the reference state for agent $i$ at time $t$. Typically, the reference states for each agent is a path of minimum cost they would follow in the absence of other agents in the environment, e.g., a straight line connecting the start and goal location. Furthermore, we denote the joint reference state of all agents at time $t$ as $\hat{x}_{t}$. We use $\mathop{\rm blkdiag}(A_{1},\ldots,A_{n})$ to denote the block-diagonal of the matrices $A_{1},\ldots,A_{n}$. Let $a_{0:\tau}:=\{a_{0},\ldots,a_{\tau}\}$ denote the collection of vectors $a_{t}$’s over the entire horizon $0\leq t\leq\tau$. We use $\{{x}_{0:\tau}^{i},{u}_{0:\tau}^{i}\}_{i=1}^{N}$ to denote joint trajectories of all $N$ agents across all times. Note that using the joint notation, we can equivalently write the joint trajectories of all agents as $\{{x}_{0:\tau},{u}_{0:\tau}\}$.
III-A Local Generalized Nash Equilibrium Trajectories
Due to the interactive nature of the problem, generally, it is not possible for all agents to minimize their costs simultaneously while satisfying the constraints. Therefore, the interaction outcome is best captured by local GNE trajectories of the underlying trajectory game. With dynamics as (1) and constraints as (2), we denote the game in the compact form by
$\mathcal{G}:=\left(\mathcal{N},\{Q^{i}\}_{i\in\mathcal{N}},\{Q^{i}_{\tau}\}_{i\in\mathcal{N}},\{R^{i}\}_{i\in\mathcal{N}},g,f,\hat{x}_{0:\tau}\right).$
For a given $\epsilon$, we define the set of local trajectories around a joint trajectory as $\mathcal{D}\left(\{{x}_{0:\tau},{u}_{0:\tau}\},\epsilon\right)=\left\{\{{x}^{\prime}_{0:\tau},{u}^{\prime}_{0:\tau}\}\bigg{|}\sum_{t=0}^{\tau}\left(\left\lVert x_{t}-{x}^{\prime}_{t}\right\rVert^{2}_{2}+\left\lVert u_{t}-{u}^{\prime}_{t}\right\rVert^{2}_{2}\right)\leq\epsilon\right\}.$ A local GNE trajectory for the game $\mathcal{G}$ is then defined as follows.
Definition 1.
A set of joint trajectories $\{{x}_{0:\tau}^{i^{\star}},{u}_{0:\tau}^{i^{\star}}\}_{i=1}^{N}$ is a local GNE trajectory for $\mathcal{G}$ if for all agent $i\in\mathcal{N}$, agent i’s trajectory $\{{x}_{0:\tau}^{i^{\star}},{u}_{0:\tau}^{i^{\star}}\}$ is an optimal solution of the following optimization problem withing a local neighborhood $\mathcal{D}\left(\{{x}^{\star}_{0:\tau},{u}^{\star}_{0:\tau}\},\epsilon\right)$ for some $\epsilon>0$:
$\begin{array}[]{ll}\underset{\{x_{0:\tau}^{i},u_{0:\tau}^{i}\}}{\text{minimize}}&\frac{1}{2}\left\lVert x_{\tau}^{i}-\hat{x}_{\tau}^{i}\right\rVert_{Q^{i}_{\tau}}^{2}+\frac{1}{2}\sum_{t=0}^{\tau-1}\left\lVert x_{t}^{i}-\hat{x}_{t}^{i}\right\rVert_{Q^{i}}^{2}\\ &\qquad\qquad+\frac{1}{2}\sum_{t=0}^{\tau}\left\lVert u_{t}^{i}\right\rVert_{R^{i}}^{2}\\ \mbox{subject to}&x_{0}^{i}=\hat{x}_{0}^{i},\enskip{x}_{t+1}^{i}=f^{i}({x}_{t}^{i},{u}_{t}^{i}),\enskip 0\leq t\leq\tau-1,\\ &g(\tilde{x}_{t},\tilde{u}_{t})\leq{0}_{c},\enskip 0\leq t\leq\tau,\\ \end{array}$ (4)
where we use the following notation in $g(\tilde{x}_{t},\tilde{u}_{t})$
$\tilde{x}_{t}$ $\coloneqq\begin{bmatrix}({x}_{t}^{1^{\star}})^{\top}&({x}_{t}^{2^{\star}})^{\top}&\cdots&({x}_{t}^{{i-1}^{\star}})^{\top}&(x_{t}^{i})^{\top}&({x}_{t}^{{i+1}^{\star}})^{\top}&\cdots&({x}_{t}^{N^{\star}})^{\top}\end{bmatrix}^{\top},$
$\tilde{u}_{t}$ $\coloneqq\begin{bmatrix}({u}_{t}^{1^{\star}})^{\top}&({u}_{t}^{2^{\star}})^{\top}&\cdots&({u}_{t}^{{i-1}^{\star}})^{\top}&(u_{t}^{i})^{\top}&({u}_{t}^{{i+1}^{\star}})^{\top}&\cdots&({u}_{t}^{N^{\star}})^{\top}\end{bmatrix}^{\top}.$
Intuitively, at a local GNE, no agent can reduce their cost function by independently changing their trajectory to any alternative feasible trajectory within a local neighborhood of the joint equilibrium trajectory, $\{{x}_{0:\tau}^{i^{\star}},{u}_{0:\tau}^{i^{\star}}\}_{i=1}^{N}$. Note that in generalized Nash equilibria, the agents’ trajectories are coupled both through their independent objective functions and their joint constraints. It is important to highlight that solving (4) requires solving a set of $N$ coupled constrained optimal control problems, which is often computationally expensive. Therefore, finding all the local Nash equilibria of the game can be time-consuming when using traditional game theoretic methods that aim to solve coupled optimal control problems. Hence, in order to alleviate this problem, we leverage constrained potential games, which we build on to identify different modes of interaction.
IV Constrained Trajectory Potential Games
Constrained trajectory potential games are a class of games for which a GNE of the original game can be computed by optimizing a single constrained optimal control problem called Potential, as described in [13]. This reduces the computational cost of computing the GNEs of the game. In the following, we will prove that game $\mathcal{G}$ can be expressed as a constrained trajectory potential game. The following theorem shows that any local minimizer of such a potential function is a local generalized Nash equilibrium trajectory.
Theorem 1.
A set of joint trajectories $\{{x}_{0:\tau}^{i^{\star}},{u}_{0:\tau}^{i^{\star}}\}_{i=1}^{N}$ is a local GNE trajectory if within a local neighborhood $\mathcal{D}\left(\{{x}^{\star}_{0:\tau},{u}^{\star}_{0:\tau}\},\epsilon\right)$, it is an optimal solution of
$\begin{array}[]{ll}\underset{\{x_{0:\tau},u_{0:\tau}\}}{\text{minimize}}&\frac{1}{2}\left\lVert x_{\tau}-\hat{x}_{\tau}\right\rVert_{Q_{\tau}}^{2}+\frac{1}{2}\sum_{t=0}^{\tau-1}\left\lVert x_{t}-\hat{x}_{t}\right\rVert_{Q}^{2}\\ &\qquad+\frac{1}{2}\sum_{t=0}^{\tau}\left\lVert u_{t}\right\rVert_{R}^{2}\\ \mbox{subject to}&x_{0}=\hat{x}_{0},\enskip x_{t+1}=f(x_{t},u_{t}),\enskip 0\leq t\leq\tau-1,\\ &g(x_{t},u_{t})\leq{0}_{c},\enskip 0\leq t\leq\tau,\end{array}$ (5)
for some $\epsilon>0$, where $Q_{\tau}\coloneqq\mathop{\rm blkdiag}(Q^{1}_{\tau},\ldots,Q^{N}_{\tau}),\enskip Q\coloneqq\mathop{\rm blkdiag}(Q^{1},\ldots,Q^{N})$ and $R\coloneqq\mathop{\rm blkdiag}(R^{1},\ldots,R^{N})$.
Proof.
This theorem follows from Theorem 2 from [13], which states that if each agent’s cost function depends only on their own states and actions, then the underlying game is a constrained dynamic potential game. It is easy to see that agent costs according to (3) depend only on the individual agent’s own states and control inputs. Therefore, we can use the results from Theorem 2 from [13] to obtain the potential function of the game $\mathcal{G}$ to be
$\sum_{i=1}^{N}(\frac{1}{2}\left\lVert x_{\tau}^{i}-\hat{x}_{\tau}^{i}\right\rVert_{Q^{i}_{\tau}}^{2}+\frac{1}{2}\sum_{t=0}^{\tau-1}\left\lVert x_{t}^{i}-\hat{x}_{t}^{i}\right\rVert_{Q^{i}}^{2}+\frac{1}{2}\sum_{t=0}^{\tau}\left\lVert u_{t}^{i}\right\rVert_{R^{i}}^{2}).$ (6)
Furthermore, using the joint notations, we can re-write the sums in (6) as $\sum_{i=1}^{N}\left(\frac{1}{2}\left\lVert x_{\tau}^{i}-\hat{x}_{\tau}^{i}\right\rVert_{Q^{i}_{\tau}}^{2}\right)=\frac{1}{2}\left\lVert x_{\tau}-\hat{x}_{\tau}\right\rVert_{Q_{\tau}}^{2},$ $\sum_{i=1}^{N}\left(\frac{1}{2}\left\lVert x_{t}^{i}-\hat{x}_{t}^{i}\right\rVert_{Q^{i}}^{2}\right)=\frac{1}{2}\left\lVert x_{t}-\hat{x}_{t}\right\rVert_{Q}^{2},$ $\sum_{i=1}^{N}\left(\frac{1}{2}\left\lVert u_{t}^{i}\right\rVert_{R^{i}}^{2}\right)=\frac{1}{2}\left\lVert u_{t}\right\rVert_{R}^{2}$. which in turn gives us (5).
Therefore, we have proven that $\mathcal{G}$ is a constrained trajectory potential game with potential (5). Hence, using Theorem 2 from [13], any local solution of (5) will also be a local GNE of the game $\mathcal{G}$. ∎
Therefore, using Theorem 1, finding local GNE trajectories is equivalent to finding local solutions to (5). The conventional approach to solving the optimization in (5) relies on techniques such as inter-point methods [23] or sequential quadratic programming [24]. Naively, the local solutions can be obtained by exhaustive initialization of random trajectories to an optimization solver, such as IPOPT [23], hoping to explore the solution space of local Nash equilibria. However, random exploration using gradient-based methods becomes computationally prohibitive for large $N$ and inefficient as the same local equilibria can be arrived at from multiple random initializations. To address this, we utilize a Bayesian inference perspective in the following section to recover all local minima of (5) in an efficient manner.
V Identifying Local Nash Equilibria via Bayesian Inference
In this section, we propose our approach to solve (5) from a Bayesian inference perspective. We reformulate the optimal control problem (5) into an equivalent Bayesian inference problem by utilizing the reference trajectory and constraint violations as virtual measurements that provide evidence for the optimal inference of the system trajectory in (5). At its core, the Bayesian inference problem is about approximating the probability density function (PDF) of the trajectories given measurement information derived from the objective and constraints in (5). Then, the high probability regions of the PDF to be approximated will correspond to the local GNEs of interest. This enables using the sample efficient implicit particle filter [30], which is designed to sample from the high probability regions of the PDF. The resulting identified equilibria then serve as a warm-start initialization for a downstream optimizer such as IPOPT, which finds the fine-tuned trajectory. We first formulate the MulitNash-PF problem and then outline the algorithm that realizes the framework.
V-A MultiNash-PF framework via Bayesian Inference
Given problem (5), we view the objective function and constraints as a source of optimality information to facilitate the inference of the optimal trajectory. This motivates the introduction of a state-space model that sets us up to construct the Bayesian reformulation as a joint state and control inference via the objective function and constraint as measurements. We construct the virtual system as
$x_{t+1}$ $=f(x_{t},u_{t}),$ $y_{x,t}$ $=x_{t}+v_{t},$
$u_{t+1}$ $=w_{t},$ $y_{g,t}$ $=\psi\left(g\left(x_{t},u_{t}\right)\right)+\eta_{t},$ (7)
for $0\leq t\leq\tau$, where $x_{t}$ and $u_{t}$ are the states of the virtual system, $y_{x,t}$, $y_{g,t}$ are measurements, and $v_{t}$, $w_{t}$, and $\eta_{t}$ are bounded disturbances. The function $\psi(\cdot)$ represents a barrier function that can capture the degree of constraint violation by a trajectory. We chose the softplus barrier function for its continuity in the forthcoming filtering algorithm. The solfplus function is expressed as:
$\psi(g\left(x_{t},u_{t}\right))=\ln\left(1+\exp(g(x_{t},u_{t}))\right)/\alpha,$ (8)
where $\alpha$ is a tuneable parameter that controls the strictness of enforcing constraints. Considering the virtual system (V-A), we relax the problem (5) using soft constraints as:
$\begin{array}[]{ll}\underset{\{v_{t},w_{t},\eta_{t}\}_{t=0}^{\tau}}{\min}&\frac{1}{2}\left\lVert v_{\tau}\right\rVert_{Q_{\tau}}^{2}+\frac{1}{2}\sum_{t=0}^{\tau-1}\left\lVert v_{t}\right\rVert_{Q}^{2}\\ &+\frac{1}{2}\sum_{t=0}^{\tau}\left\lVert w_{t}\right\rVert_{R}^{2}+\frac{1}{2}\sum_{t=0}^{\tau}\left\lVert\eta_{t}\right\rVert^{2}_{Q_{\eta}},\\ \quad\;\mathrm{s.t.}&\begin{bmatrix}x_{0}\\ u_{0}\end{bmatrix}=\begin{bmatrix}\hat{x}_{0}\\ w_{-1}\end{bmatrix},\begin{bmatrix}x_{t+1}\\ u_{t+1}\end{bmatrix}=\begin{bmatrix}f(x_{t},u_{t})\\ 0_{m}\end{bmatrix}+\begin{bmatrix}0_{n}\\ w_{t}\end{bmatrix},\vspace{0.2cm}\\ &\begin{bmatrix}y_{x,t}\\ y_{g,t}\end{bmatrix}=\begin{bmatrix}x_{t}\\ \phi\left(g(x_{t},u_{t})\right)\end{bmatrix}+\begin{bmatrix}v_{t}\\ \eta_{t}\end{bmatrix},0\leq t\leq\tau,\end{array}$ (9)
where $Q_{\eta}$ is a positive definite weight matrix that assigns the importance of constraint satisfaction. Problem (9) differs from (5) in that it considers a soft inequality constraint using the barrier function (8). Further, (9) is also known as the moving horizon estimation problem [31] which aims to find the minima of the disturbances $\left\{v_{t}^{\star},w_{t}^{\star},\eta_{t}^{\star}\right\}_{t=0}^{\tau}$. A direct attempt at solving (9) using traditional gradient-based methods poses a similar computational burden in identifying local Nash equilibria compared to problem (5). However, the above connection between (5) and the moving horizon estimation problem (9) highlights the potential to treat the problem from a Bayesian inference perspective. It can be seen that the constraints in (9) are the same as the considered virtual system in (V-A), suggesting that state inference can be performed via efficient estimation methods. To proceed, we rewrite (V-A) compactly as
$\bar{x}_{t+1}=\bar{f}(\bar{x}_{t})+\bar{w}_{t},\;{\bar{y}}_{t}=\bar{h}(\bar{x}_{t})+\bar{v}_{t},$ (10)
where
$\bar{x}_{t}$ $=\begin{bmatrix}x_{t}\\ u_{t}\end{bmatrix},\>\bar{y}_{t}=\begin{bmatrix}y_{x,t}\\ y_{g,t}\end{bmatrix},\>\bar{w}_{t}=\begin{bmatrix}0_{n}\\ w_{t}\end{bmatrix},$
$\bar{v}_{t}=\begin{bmatrix}v_{t}\\ \eta_{t}\end{bmatrix},\;\bar{f}($ $\bar{x}_{t})=\begin{bmatrix}f(x_{t},u_{t})\\ {0}_{m}\end{bmatrix},\;\bar{h}(\bar{x}_{t})=\begin{bmatrix}x_{t}\\ \psi(g\left(x_{t},u_{t}\right))\end{bmatrix}.$
Here, we relax $\bar{w}_{t}$ and $\bar{v}_{t}$ for $0\leq t\leq\tau$ to be stochastic disturbances modeled as normally distributed random variables with PDF
$\bar{v}_{t}\sim\mathcal{N}(0_{n+c},\bar{Q}),\;\bar{v}_{\tau}\sim\mathcal{N}(0_{n+c},\bar{Q}_{\tau}),\;\bar{w}_{t}\sim\mathcal{N}(0_{n+m},\bar{R}),$ (11)
with $\bar{Q}\coloneqq\mathop{\rm blkdiag}(Q^{-1},Q_{\eta}^{-1})$, $\bar{Q}_{\tau}\coloneqq\mathop{\rm blkdiag}(Q_{\tau}^{-1},Q_{\eta}^{-1})$, $\bar{R}\coloneqq\mathop{\rm blkdiag}(0_{n\times n},R^{-1})$. The state estimation relevant to (10) is to estimate the state $\bar{x}_{t}$ given the measurements $\bar{y}_{t}$. We must set the virtual measurement’s value to correspond to the objective of the optimization in (5). We achieve this by noting that $y_{x,t}$ is a measurement value used to correct the state estimate of $x_{t}$ towards the reference $\hat{x}_{t}$. While $y_{g,t}$ is to satisfy constraints through (8). Further, we observe that the barrier function in (8) outputs zero when the constraints are satisfied. Hence, the virtual measurement that will drive $\bar{x}_{t}$ towards optimality takes the value
$\bar{y}_{t}=\begin{bmatrix}\hat{x}_{t}\\ {0}_{c}\end{bmatrix},0\leq t\leq\tau.$
The Bayesian inference problem pertaining to (10) is to characterize the conditional PDF $p(\bar{x}_{0:\tau}\,|\,\bar{y}_{0:\tau})$. In the following, lemma (1) elucidates that characterizing the maximum a posteriori estimate of $p(\bar{x}_{0:\tau}\,|\,\bar{y}_{0:\tau})$ holds equivalence to solving (5).
Lemma 1.
Assuming independence between the noise vectors $\bar{v}_{t}$ and $\bar{w}_{t}$ in (11), the problem in (5) has the same optima as
$\bar{x}_{0:\tau}^{\star}=\arg\max_{\bar{x}_{0:\tau}}\>\log p(\bar{x}_{0:\tau}\,|\,\bar{y}_{0\tau}).$
We refer the interested reader to [12, Thm. 1] for the proof. The above lemma implies that the local GNEs in (5) translate to multiple local maximums of $p(\bar{x}_{0:\tau}\,|\,\bar{y}_{0:\tau})$. Hence, the Bayesian inference problem for (10) involves characterizing the multi-modal PDF $p(\bar{x}_{0:\tau}\,|\,\bar{y}_{0:\tau})$. Generally, state estimation of a multi-modal PDF does not admit closed-form solutions [32]. Hence, we reside to approximate solutions. A powerful approach is to use Monte Carlo methods that empirically approximate the distribution via a set of $J$ particles as
$p(\bar{x}_{0:\tau}\,|\,\bar{y}_{0:\tau})\approx\sum_{j=1}^{J}w_{\tau}^{j}\delta\left(x_{0:\tau}-x_{0:\tau}^{j}\right),$ (12)
where $\delta(\cdot)$ is the dirac delta function and $w_{\tau}^{i}$ is the weight assigned to a sample trajectory $\bar{x}_{0:\tau}^{i}$ for $j=1,\ldots,J$. After identifying the local GNE through the above approximation, we extract the trajectories that appear as local GNE and use them to initialize an optimization solver. This completes the exposition of the MultiNash-PF framework.
We employ the implicit particle filter  [30] due to its two-fold benefit for implementing the MultiNash-PF framework. First, as suggested in lemma 1, the local maximums (i.e high-probability regions) of $p(\bar{x}_{0:\tau}\,|\,\bar{y}_{0:\tau})$ correspond to the local GNEs. Hence, focusing our sampling efforts towards these high-probability regions can reduce the required samples for the approximation in (12). Second, we want to avoid the traditionally needed resampling step in particle filtering methods to ensure that the inferred trajectories are smooth. Fortunately, the implicit particle filter satisfies both of the aforementioned criteria. We leverage the Markovian property of the virtual system (10) to approximate (12). Using the Bayes’ rule, we have
$p(\bar{x}_{0:t}\,|\,\bar{y}_{0:t})=p(\bar{y}_{t}\,|\,\bar{x}_{t})p(\bar{x}_{t}\,|\,\bar{x}_{t-1})p(\bar{x}_{0:t-1}\,|\,\bar{y}_{0:t-1}),$ (13)
for $0\leq t\leq\tau$. This indicates that given the particle $\bar{x}_{t-1}^{j}$ from the prior distribution $p(\bar{x}_{0:t-1}\,|\,\bar{y}_{0:t-1})$, we can recursively obtain $\bar{x}_{t}^{j}$, reducing the complexity involved in directly sampling from a high-dimensional PDF. As shown in [30], we use implicit importance sampling and identify an implicit mapping to sample from high probability regions of (13). For this purpose, we make a local Gaussian assumption around $\bar{x}_{t-1}^{j}$ such that
$p(\bar{x}_{t}^{j}|\bar{y}_{0:t})\sim\mathcal{N}\left(\hat{m}_{t}^{j},\hat{\Sigma}_{t}^{x,j}\right),$ (14)
where $\hat{m}_{t}^{j}$ and $\hat{\Sigma}_{t}^{x,j}$ are the mean and error covariance of $\bar{x}_{t}^{j}$, respectively. These quantities are then found recursively by using a nonlinear Kalman filter. Then, the implicit map $X(\gamma_{t}^{j})$ can be computed in closed form as
$\bar{x}_{t}^{j}=\hat{m}_{t}^{j}+\sqrt{\hat{\Sigma}_{t}^{x,j}}\gamma_{t}^{j},$ (15)
where $\gamma_{t}^{j}$ is a sample from $\mathcal{N}(0,I_{n})$ and $I_{n}$ denotes the identity matrix of size $n$.
Algorithm 1 MultiNash-PF: Multiple local GNEs based on particle filter
1:Formulate the constrained dynamic potential game (5)
2:Setup the virtual model (10)
3:Perform unscented implicit particle filtering as follows:
4:Initialize the particles: $\bar{x}_{0}^{j}$, $\hat{\Sigma}^{x,j}_{0}$, and $w_{0}^{j}$ for $j=1,\ldots,J$
5:for $t=0,1,\ldots,\tau-1$ do
6:     for $i=1,2,\ldots,J$ do
7:Run Kalman prediction step:
8:         $\left(m_{t+1}^{j},\Sigma^{x,j}_{t+1}\right)\leftarrow\text{{\bf UT}}(\bar{x}_{t}^{j},\Sigma_{t}^{x,j},\bar{R},\bar{f})$
9:Run Kalman update step:
10:         if $t\leq\tau-2$ then
11:              $(\hat{\bar{y}}_{t+1}^{j},\Sigma^{y,j}_{t+1},\Sigma^{xy,j}_{t+1})\leftarrow\text{{\bf UT}}(m_{t+1}^{j},\Sigma_{t+1}^{x,j},\bar{Q},\bar{h})$
12:         else
13:              $(\hat{\bar{y}}_{t+1}^{j},\Sigma^{y,j}_{t+1},\Sigma^{xy,j}_{t+1})\leftarrow\text{{\bf UT}}(m_{t+1}^{j},\Sigma_{t+1}^{x,j},\bar{Q}_{\tau},\bar{h})$
14:         end if
15:         $K_{t+1}^{j}\leftarrow\Sigma_{t+1}^{xy,j}(\Sigma_{t+1}^{y,j})^{-1}$
16:         $\hat{m}^{x,j}_{t+1}=m_{t+1}^{j}+K_{t+1}^{j}\left(\bar{y}_{t+1}-\hat{\bar{y}}_{t+1}^{j}\right)$
17:         $\hat{\Sigma}_{t+1}^{x,j}\leftarrow\Sigma_{t+1}^{x,j}-K_{t+1}^{j}\Sigma^{y,j}_{t+1}(K_{t+1}^{j})^{\top}$
18:Draw a high-probability sample $\gamma_{t+1}^{j}\sim\mathcal{N}(0,I_{n})$ for the implicit mapping in (15):
19:         $\bar{x}_{t+1}^{j}\leftarrow\hat{m}_{t+1}^{x,j}+\sqrt{\hat{\Sigma}_{t+1}^{x,j}}\gamma_{t+1}^{j}$
20:Weight and normalize samples:
21:         $w_{t+1}^{j}=\frac{w_{t}^{j}p(\bar{y}_{t+1}\,|\,\bar{x}_{t}^{j})}{\sum_{j=1}^{J}w_{t}^{j}p(\bar{y}_{t+1}\,|\,\bar{x}_{t}^{j})},\qquad p(\bar{y}_{t+1}\,|\,\bar{x}_{t}^{j})\sim\mathcal{N}(\hat{\bar{y}}_{t+1}^{j},\Sigma^{y,j}_{t+1})$
22:     end for
23:     Perform resampling if necessary
24:end for
25:Extract trajectories from $\{x_{0:\tau}\}_{j=1}^{J}$ and warm start IPOPT solver to obtain the local GNEs $\{{x}_{0:\tau}^{i^{\star}},{u}_{0:\tau}^{i^{\star}}\}_{i=1}^{N}$ of $\mathcal{G}$.
V-B Implicit Particle Filter via Unscented Kalman filter
As in [30], we note that the local Gaussian approximation in (14) can be recursively computed through a nonlinear Kalman filter evolving according to (13) for all $j=1,\ldots,J$. Our nonlinear Kalman filter of choice is the unscented Kalman filter due to its high filtering accuracy for low-dimensional state space models [33]. The unscented Kalman filter is based on the unscented transform, which utilizes deterministic samples, known as sigma points, to trace the evolution of the mean and covariance of a Gaussian random variable through a nonlinear function. The unscented Kalman filter [33] utilizes the unscented transform in the Kalman prediction and measurement update steps to keep track of the Gaussian statistics through the state transition and measurement function in (10). From now on, we use UT to denote Unscented Transform algorithm.
Using the unscented Kalman filter, we can recursively compute the local Gaussian approximation (14) at time $t$. Then, we implicitly map the particle to the high-probability regions by sampling $\gamma_{t}^{i}\sim\mathcal{N}(0,I_{n})$ for (15) to obtain the particle $\bar{x}_{t}^{i}$. If necessary, we perform resampling for the particle trajectory $\bar{x}_{0:t}^{i}$. This process is repeated until $t=\tau$ for each particle to obtain the set of trajectories $\{\bar{x}_{0:\tau}\}_{j=1}^{J}$, which will include the coarse estimates of local GNEs. This completes the unscented implicit particle filter, which has identified the coarse estimates of multiple local Nash equilibria in (5). As the last step, the estimates of local equilibria are extracted and used as warm start initialization to the IPOPT optimizer to find the optimal local GNEs. To extract the equilibria, we carefully observe the trajectories and select the trajectories that are repeatedly taken by multiple particles. The pragmatic extraction of equilibria from the particle trajectories is nontrivial and is left for future study. The implementation details of the proposed MultiNash-PF through the unscented implicit particle filter are provided in Algorithm 1.
VI Numerical Simulations
To showcase the capabilities of our method, we take motivation from real-life highway scenarios. We first consider an interaction between two agents moving toward each other as shown in Fig 1. We consider the following discrete-time unicycle dynamics to model each agent $i\in\{1,2\}$:
$p^{i}_{t+1}$ $=p^{i}_{t}+\Delta t\cdot v^{i}_{t}\cos{\theta^{i}_{t}},\;q^{i}_{t+1}=q^{i}_{t}+\Delta t\cdot v^{i}_{t}\sin{\theta^{i}_{t}}$
$\theta^{i}_{t+1}$ $=\theta^{i}_{t}+\Delta t\cdot\omega_{i},\;v^{i}_{t+1}=v^{i}_{t}+\Delta v^{i}_{t},\;\omega^{i}_{t+1}=\omega^{i}_{t}+\Delta\omega^{i}_{t},$
where $\Delta t$ is the time-step size, $p^{i}_{t}$ and $q^{i}_{t}$ are $x$ and $y$ coordinates of the positions, $\theta^{i}_{t}$ is the heading angle from positive x-axis, $v^{i}_{t}$ is the linear velocity, and $\omega^{i}_{t}$ is the angular velocity of the agent $i$ at time $t$. Further, $\Delta v^{i}_{t}$ is the change in linear velocity, and $\Delta\omega^{i}_{t}$ is the change in angular velocity of the agent $i$ at time $t$, respectively. For each agent $i$, we denote the state vector $x^{i}_{t}$ and the action vector $u^{i}_{t}$ as follows
$x^{i}_{t}=\begin{bmatrix}p^{i}_{t}&q^{i}_{t}&\theta^{i}_{t}&v^{i}_{t}&\omega^{i}_{t}\end{bmatrix}^{\top},\enskip u^{i}_{t}=\begin{bmatrix}\Delta v^{i}_{t}&\Delta\omega^{i}_{t}\end{bmatrix}^{\top}.$
We formulate this scenario as a two-player game according to cost functions as described in (3). The reference trajectory for both agents is given as a straight line on the highway, starting from their initial position to their respective goal location. With $\hat{Q}=diag(50,10,5,5,2)$, the reference tracking and control effort cost matrices are set as
$Q^{i}=0.6\hat{Q},\;Q^{i}_{\tau}=100\hat{Q},\;R^{i}=diag(8,4).$ (16)
for $i\in\{1,2\}$. In addition, the agents are subject to the constraint of avoiding collision with each other, which we define as
$g_{c}(x_{t},u_{t})\coloneq-d(x^{1}_{t},x^{2}_{t})+r_{\text{col}}\leq 0,$ (17)
where $d(x^{1}_{t},x^{2}_{t})\coloneq\sqrt{(p^{1}_{t}-p^{2}_{t})^{2}+(q^{1}_{t}-q^{2}_{t})^{2}}$ is the Euclidean distance between the two agents and $r_{\text{col}}\coloneq 3m$ is the collision avoidance radius between the two agents.
With this setup, we employ MultiNash-PF from Algorithm 1 to efficiently capture the multi-modality in the solution space and then extract the solutions to provide a warm start trajectory to a solver like IPOPT [23] to recover different local GNEs. We use the Julia programming language and utilize the JuMP [34] library to solve the game using IPOPT solver. The results of our method are provided in Fig 1. As we can see, the presence of two agents and collision avoidance constraints induces two different possible Nash equilibria. Either both the agents will yield to their right, or they will yield to their left to avoid collision. As shown in Fig 1, our method can recover both equilibria.
Furthermore, we consider a more complicated scenario in which two agents aim to exchange their positions while avoiding an obstacle of radius $r_{\text{obs}}=4m$ located at the midpoint of the two agents. It should be noted that the radius of the obstacle is larger than the collision avoidance radius of the two agents.
We consider the same unicycle dynamics as before, and the reference trajectory cost is also the same as in (16). Furthermore, the collision avoidance radius between the two agents is also the same as before. In addition to (17), the constraints for this game will include both agents avoiding the obstacle $g^{i}_{o}(x_{t},u_{t}):=\begin{bmatrix}-d(x^{i}_{t},x_{\text{obs}})+r_{\text{obs}}\end{bmatrix}$ for $i=1,2$ where $x_{\text{obs}}$ is the location of the obstacle. We also assume that we have constraints on the control bounds $g_{b}(x_{t},u_{t})=\begin{bmatrix}-v^{1}_{t},&-v^{2}_{t},&|u^{1}_{t}|-u^{1}_{b},&|u^{2}_{t}|-u^{2}_{b}\end{bmatrix}$, where $u^{1}_{b}$ and $u^{2}_{b}$ are controlled action bounds for agents 1 and 2, respectively. We write the combined constraints as $g(x_{t},u_{t}):=\begin{bmatrix}g_{c}&g^{1}_{o}&g^{2}_{o}&g_{b}\end{bmatrix}^{\top}\leq 0.$ We choose the values of control action bounds to be $u^{1}_{b}=u^{2}_{b}=[0.15,0.75]^{\top}$. This setup induces multitudes of local equilibria. Two trivial local equilibria are when the two agents pass each other through opposite sides of the obstacle. However, there exist four other nontrivial equilibria when both agents can go on the same side of the obstacle, and one of the agents can yield to the other one. It is challenging to recover all six equilibria without utilizing the capabilities of the implicit particle filter. Upon performing simulation experiments with our method, we observe that MultiNash-PF is able to recover all six Nash equilibria as shown in Fig. 2.
(a) Histogram of the solve time comparison of MultiNash-PF and the baseline.
(b) Histogram of the number of the IPOPT runs required to obtain all six equilibria.
Figure 3: Histograms of the comparison of MultiNash-PF with the baseline for the experiment shown in Fig. 2.
(a) The average solve time for MultiNash-PF is $3.36\pm 0.59s$ while the baseline’s average solve time is $6.71\pm 2.83s$. MultiNash-PF reduces the solve time by about $50\%$ compared to the baseline while giving a much lower variance on the solve time. (b) For the baseline, the average number of IPOPT runs required to obtain all six different equilibria is $18.64\pm 7.57$, while for MultiNash-PF, this is always $6$. MultiNash-PF requires 3 times fewer IPOPT runs.
VI-1 Comparison with Baseline
In order to compare the effectiveness of our method, we consider a baseline method. In the baseline method, we directly initialize the IPOPT solver with random perturbations of reference trajectories and compute the resulting optimal solutions. For our method, we fixed $J=50$ particles and obtained the solution trajectories using implicit particle filtering from Algorithm 1 described in Section V. After that, we extract all the different equilibria and provide them as a warm starting point for IPOPT to obtain the final solutions. The compute time of our method is a combination of time used in implicit particle filtering and time used by IPOPT to obtain all the different equilibria given warm starting from filtering.
We run the IPOPT solver until we recover all 6 different equilibria. We repeat this experiment for 100 different Monte Carlo runs. We observe that it requires, on average, $18.64\pm 7.57$ number of IPOPT runs to obtain all six different equilibria using the baseline approach of random warm-starting. However, with our method, the number of IPOPT runs required is always six because we always obtain all six modes of equilibria from implicit particle filtering. A histogram of the number of IPOPT runs required for obtaining all six equilibria is plotted in Fig. 3(b).
Furthermore, the solve time to recover all 6 different Nash equilibria for the baseline is $6.71\pm 2.83s$, while for MultiNash-PF, the solve time is $3.36\pm 0.59s$. Our method results in a solve-time reduction of $50\%$ compared to the baseline while giving a much lower variance on the solve time. A histogram of the solve times of the baseline compared to our method is plotted in Fig. 3(a). As explained earlier, our method’s solve time is a combination of the time used in implicit particle filtering and the time used by IPOPT. We observe that the time required for implicit particle filtering over 100 Monte Carlo runs is $0.37\pm 0.072s$ while the time used by IPOPT to obtain the final solutions is $2.99\pm 0.57s$. Therefore, in our MultiNash-PF, the time used by implicit particle filtering is $11\%$ of the total solve time, which shows that filtering is very efficient in exploring the solution space of local Nash equilibria and contributes really less to the overall computation time of MultiNash-PF.
VII Conclusion
In this work, we introduced MultiNash-PF, a novel algorithm designed to efficiently compute multiple local generalized Nash equilibria (GNEs) in constrained trajectory games. By integrating potential game theory with implicit particle filtering, MultiNash-PF provides coarse estimates of local minimizers, which are then refined using off-the-shelf optimization solvers. Our experimental results in two-player trajectory games involving nonlinear dynamics and collision avoidance constraints demonstrate that MultiNash-PF not only identifies all local Nash equilibria but also significantly reduces computation time compared to traditional random initialization methods. This enhanced efficiency and accuracy underline the potential of MultiNash-PF to improve identifying various interaction modes in complex, real-world environments, paving the way for more robust and socially acceptable autonomous systems.

### v2

@@
 Abstract
-Modern-world robotics involves complex environments where multiple autonomous agents must interact with each other and other humans.
-This necessitates advanced interactive multi-agent motion planning techniques.
-Generalized Nash equilibrium(GNE), a solution concept in constrained game theory, provides a mathematical model to predict the outcome of interactive motion planning, where each agent needs to account for other agents in the environment.
-However, in practice, multiple local GNEs may exist.
-Finding a single GNE itself is complex as it requires solving coupled constrained optimal control problems.
-Furthermore, finding all such local GNEs requires exploring the solution space of GNEs, which is a challenging task.
-This work proposes the MultiNash-PF framework to efficiently compute multiple local GNEs in constrained trajectory games.
-Potential games are a class of games for which a local GNE of a trajectory game can be found by solving a single constrained optimal control problem.
-We propose MultiNash-PF that integrates the potential game approach with implicit particle filtering, a sample-efficient method for non-convex trajectory optimization.
-We first formulate the underlying game as a constrained potential game and then utilize the implicit particle filtering to identify the coarse estimates of multiple local minimizers of the game’s potential function.
+Modern robotic systems frequently engage in complex multi-agent interactions, many of which are inherently multi-modal, meaning they can lead to multiple distinct outcomes.
+To interact effectively, robots must recognize the possible interaction modes and adapt to the one preferred by other agents.
+In this work, we propose an efficient algorithm for capturing the multimodality in multi-agent interactions.
+We leverage a game-theoretic planner to model interaction outcomes as equilibria where each equilibrium corresponds to a distinct interaction mode.
+We then develop an efficient algorithm to identify all the equilibria, allowing robots to reason about multiple interaction modes.
+More specifically, we formulate interactive planning as Constrained Potential Trajectory Games (CPTGs) and model interaction outcomes by local Generalized Nash equilibria (GNEs) of the game.
+CPTGs are a class of games for which a local GNE can be found by solving a single constrained optimal control problem where a potential function is minimized.
+We propose to integrate the potential game approach with implicit particle filtering, a sample-efficient method for non-convex trajectory optimization.
+We utilize implicit particle filtering to identify the coarse estimates of multiple local minimizers of the game’s potential function.
 MultiNash-PF then refines these estimates with optimization solvers, obtaining different local GNEs.
-We show through numerical simulations that MultiNash-PF reduces computation time by up to 50% compared to a baseline approach.
+We show through numerical simulations that MultiNash-PF reduces computation time by up to 50% compared to a baseline.
+We further demonstrate the effectiveness of our algorithm in real-world human-robot interaction scenarios, where it successfully accounts for the multi-modal nature of interactions and resolves potential conflicts in real-time.
 I Introduction
@@
 For instance, autonomous cars on highways need to account for other vehicles while planning their motion.
-Many robotics applications require motion planning for selfish agents sharing a common operational space, such as autonomous cars and delivery robots.
-In such settings, autonomous agents should account for other agents in the shared environment while achieving their individual goals.
-Autonomous agents need to reason about the impact of their actions on other agents.
-Game theory provides a formal way to deal with such interactive scenarios [1].
-Furthermore, recent works have shown that game-theoretic trajectory planning makes robots more socially acceptable to humans [2].
-To model the outcome of such interactive scenarios, Nash equilibrium is a key concept in game theory.
-When multiple agents- such as robots and humans- interact, a Nash equilibrium corresponds to a joint action such that no agent has incentives to deviate from its equilibrium actions unilaterally.
-This allows us to account for agents’ individual preferences when interacting in a common environment.
-We formulate a multi-agent interactive trajectory planning problem as a trajectory game.
-In robotics, seeking Nash equilibria of the underlying trajectory games has been proven to be effective [3, 4].
-Furthermore, in real-world applications, agents may have to satisfy some safety constraints, such as avoiding collisions with obstacles and other agents in the environment.
-In such constrained multi-agent interactive settings, the interaction outcomes are captured through generalized Nash equilibrium(GNE), which are constrained Nash equilibria.
-Several recent works have shown great success in utilizing GNE for multi-agent motion planning in interactive domains [5, 6].
-However, in practice, multiple local GNEs may exist for a trajectory game.
-For example, in the two-player trajectory game illustrated in Fig.
+Most of the real-world interactions are multi-modal.
+For example, in the two-player interaction illustrated in Fig.
 1, the agents need to move towards their goals while avoiding collisions.
-Here, two local GNEs exist: clockwise trajectories and counterclockwise trajectories.
-If the two agents were to act independently, they could reach different GNEs, which would result in a collision.
-Therefore, agents need a coordination mechanism to avoid such scenarios.
-However, identifying all the possible local GNEs is required to enable effective coordination.
-Therefore, computing multiple local GNEs is critical to aligning players’ preferences over different equilibria [7] or helping coordinate players’ actions via recommendation [8].
-Furthermore, different local GNEs correspond to different modes of interactions in multi-agent interactive settings.
-Therefore, it is essential to efficiently recover all the possible multi-modal outcomes of multi-agent interactions.
+Here, two outcomes exist: both agents yielding to their right or left.
+If the two agents were to act independently, they could reach different outcomes, which would result in a collision.
+What trajectory each agent would prefer to pick depends on the convention they might be following.
+For example, in some cultures, in such settings, the convention is to yield to the right to avoid collisions while in other cultures, the convention is to yield to the left.
+Such conventions usually arise out of recurring coordination problems [1].
+Therefore, autonomous agents need to be aware of the existence of multi-modality in the interactions for efficient coordination with people who follow different conventions.
+Some recent works have shown that computing such multiple trajectories is critical to aligning agents’ preferences over different outcomes [2] or helping coordinate agents’ actions via recommendation [3].
+Figure 1: Snapshot of an interaction between a robot and a human when they exchange positions.
+Using MultiNash-PF the robot can compute both modes of interaction.
+At $t=2$, the robot is uncertain about the preferred mode of the human since the human trajectory’s distance from both equilibria is almost the same.
+At $t=4$, the human trajectory is closer to the second GNE therefore, the robot chooses to follow the second GNE accordingly and avoid collision with the human.
+To formalize this problem, we will use dynamic game theory [4] which has shown to be powerful in developing interactive and socially complaints robots [5, 6, 7, 8].
+We formulate a multi-agent interactive trajectory planning problem as a Constrained Potential Trajectory Game (CPTG).
+In CPTGs, the interaction outcomes are captured through generalized Nash equilibrium (GNE), which are constrained Nash equilibria.
+When multiple agents- such as robots and humans- interact, a GNE corresponds to a joint action such that no agent has incentives to deviate from its equilibrium actions unilaterally.
+Several recent works have shown great success in utilizing GNE for multi-agent motion planning in interactive domains [9, 10, 11, 12].
+A key challenge in interactive domains, however, is the existence of multiple equilibria, many of which are equally valid.
+We argue that each equilibrium corresponds to a distinct mode of interaction.
+For instance, in Fig.
+1, yielding to the right or left are both acceptable equilibria, representing different Nash equilibria.
+Consequently, we propose that in order to capture different interaction modes, robots must reason about multiple interaction equilibria and select the one that aligns with the convention adopted by other agents.
 Computing different local GNEs requires an efficient way to solve games.
@@
 Furthermore, computing all the different local GNEs requires initializing the game solvers with many different initial guesses of the equilibrium trajectories from prior knowledge of the specific scenario.
-Further, as the number of participating agents increases, the dimension of the nonconvex optimization grows to interactable levels.
 This further increases the required number of initial guesses for the game solvers to find all local GNEs.
@@
 First, we use the implicit particle filter to discover the coarse estimates of two different equilibria, and then utilize them as a warm starting for IPOPT to obtain the two Nash equilibrium trajectories.
-We propose a novel framework named MultiNash-PF to efficiently compute multiple local GNEs in constrained trajectory games. The approach combines
-  1. 1.
-    constrained potential games [11], a class of games for which we can compute local Nash equilibria by optimizing a single constrained optimal control problem, and,
-  2. 2.
-    constraint aware implicit particle filtering, which is a sample-efficient particle filtering method recently used for nonconvex trajectory optimization [12].
-Using constrained potential games, it is possible to compute a local GNE by solving a constrained optimal control problem using off-the-shelf optimization solvers.
-These solvers outperform state-of-the-art constrained dynamic game solvers in computation speed during [13].
-After formulating the constrained trajectory game as a constrained potential game, our approach employs the constraint-aware implicit particle filtering method to identify multiple local minima of the potential game.
+In this paper, we tackle this challenge by introducing MultiNash-PF, a novel framework designed to efficiently compute multiple local generalized Nash equilibria (GNEs) in constrained trajectory games.
+This, in turn, enables robots to recognize and reason about multiple interaction modes.
+Our approach integrates constrained potential games [15], a class of games that allow for the computation of local GNEs by solving a single constrained optimal control problem, with constraint-aware implicit particle filtering—a sample-efficient method recently applied to nonconvex trajectory optimization [16] to achieve efficient identification of multiple equilibria.
+Using CPTGs, we can compute a local GNE by solving a single constrained optimal control problem using off-the-shelf optimization solvers.
+Then, we use a constraint-aware implicit particle filtering method to identify multiple local minima of the potential minimization problem.
 This method eliminates the requirements of random initialization of the optimization solver by efficiently obtaining coarse estimates of different local GNEs.
-Next, we use the identified minima as initialization for the optimization solver.
+Next, we use the identified minima as initialization for an optimization solver to refine our solution.
 The solver then computes all the different local GNEs of the original game.
-The framework of MultiNash-PF is depicted in Fig.
-1.
 Leveraging the sample efficiency of the implicit particle filtering method, MultiNash-PF can explore and identify the multiple local GNEs faster than the random initialization of optimization solvers.
+We further show the real-time application of MultiNashPF in a human-robot navigation setting where the robot can reason about the existence of different local GNEs and plan its motion autonomously based on the modality the human prefers.
 Figure 3: MultiNash-PF identifies all the local GNE trajectories for a challenging two agent trajectory game where two agents swapped positions while avoiding an obstacle.
@@
 MultiNash-PF effectively obtains coarse estimates of local GNEs through implicit particle filtering, which are then refined using IPOPT to obtain local GNEs.
+Figure 4: When using MultiNash-PF on a difficult two agent trajectory game, we first employ an implicit particle filter to obtain coarse estimates of GNEs.
 II Related Works
@@
 Game theoretic motion planners that seek Stackelberg equilibria were proposed in [18, 7].
-Later, methods that seek Nash equilibria were developed.
-These methods differ mainly in the way they compute Nash equilibria.
 Iterative best response methods for computing Nash equilibria were developed in [19, 20].
@@
 In [2], the authors recognize the existence of multiple equilibria and focus on inferring the equilibrium of other agents in the environment.
-However, the methodology relies on computing the different approximate local Nash by rollouts of randomly selected initial strategies, which we show in this work to be computationally intractable with the risk of not identifying all possible equilibria.
+However, the methodology relies on computing the different approximate local Nash by rollouts of randomly selected initial strategies, which we show in this work performs poorly compared to our method.
 Multiple Local Solutions to Constrained Optimal Control.
@@
 Recently, in [31], an efficient particle filtering method was proposed based on the principle of implicit sampling.
-This sampling approach relies on the premise that a few particles are required for accurate estimation if the particles are sampled from high-probability regions of the probability distribution function.
-In addition, the implicit sampling technique can also mitigate the need for a resampling step to avoid filter collapse.
-III Multi-agent Constrained Trajectory Games
-We formulate the multi-agent interactive trajectory planning problem as an $N$-agent constrained trajectory game.
+III Constrained Potential Trajectory Games
+We formulate the multi-agent interactive trajectory planning problem as an $N$-agent constrained potential trajectory game (CPTG).
 Let $\mathcal{N}:=\{1,2,\ldots,N\}$ denote the set of agents’ indices.
@@
 Therefore, $f:\mathbb{R}^{n}\times\mathbb{R}^{m}\to\mathbb{R}^{n}$ will be a vector-valued dynamics function for the joint system.
-In constrained trajectory games, each agent has to satisfy some constraints that couple different agents’ states and controls and, in general, depend on other agents’ states and actions.
+In CPTGs, each agent has to satisfy some constraints that couple different agents’ states and controls.
 Let $g:\mathbb{R}^{n}\times\mathbb{R}^{m}\to\mathbb{R}^{c}$ with $c\in\mathbb{N}$ denote a continuously differentiable function that defines the constraints on the joint state $x_{t}$ and joint control $u_{t}$ at each time:
@@
 where ${0}_{c}$ is a vector of zeros in $\mathbb{R}^{c}$.
-The aim of the constraint function $g$ is to capture individual agents’ strict preferences, such as constraints for each agent avoiding obstacles, or avoiding collision with other agents.
+The aim of the constraint function $g$ is to capture individual agents’ strict preferences, such as each agent avoiding obstacles, or avoiding collision with other agents.
+It should be noted that coupling amongst agents’ decisions is captured through constraints in CPTGs.
 We assume that each agent $i$ aims to minimize an objective function.
@@
 We assume that each agent $i$ seeks to minimize the following objective:
-$\frac{1}{2}\left\lVert x_{\tau}^{i}-\hat{x}_{\tau}^{i}\right\rVert_{Q^{i}_{\tau}}^{2}+\frac{1}{2}\sum_{t=0}^{\tau-1}\left\lVert x_{t}^{i}-\hat{x}_{t}^{i}\right\rVert_{Q^{i}}^{2}+\frac{1}{2}\sum_{t=0}^{\tau}\left\lVert u_{t}^{i}\right\rVert_{R^{i}}^{2}$ (3)
+$\left\lVert x_{\tau}^{i}-\hat{x}_{\tau}^{i}\right\rVert_{Q^{i}_{\tau}}^{2}+\sum_{t=0}^{\tau-1}\left\lVert x_{t}^{i}-\hat{x}_{t}^{i}\right\rVert_{Q^{i}}^{2}+\sum_{t=0}^{\tau}\left\lVert u_{t}^{i}\right\rVert_{R^{i}}^{2}$ (3)
 where $\hat{x}_{t}^{i}\in\mathbb{R}^{n_{i}}$ is the reference state for agent $i$ at time $t$.
@@
 Therefore, the interaction outcome is best captured by local GNE trajectories of the underlying trajectory game.
-With dynamics as (1) and constraints as (2), we denote the game in the compact form by
+With dynamics as (1) and constraints as (2), we denote the game by
 $\mathcal{G}:=\left(\mathcal{N},\{Q^{i}\}_{i\in\mathcal{N}},\{Q^{i}_{\tau}\}_{i\in\mathcal{N}},\{R^{i}\}_{i\in\mathcal{N}},g,f,\hat{x}_{0:\tau}\right).$
-For a given $\epsilon$, we define the set of local trajectories around a joint trajectory as $\mathcal{D}\left(\{{x}_{0:\tau},{u}_{0:\tau}\},\epsilon\right)=\left\{\{{x}^{\prime}_{0:\tau},{u}^{\prime}_{0:\tau}\}\bigg{|}\sum_{t=0}^{\tau}\left(\left\lVert x_{t}-{x}^{\prime}_{t}\right\rVert^{2}_{2}+\left\lVert u_{t}-{u}^{\prime}_{t}\right\rVert^{2}_{2}\right)\leq\epsilon\right\}.$ A local GNE trajectory for the game $\mathcal{G}$ is then defined as follows.
+For a given $\epsilon$, we define the set of local trajectories around a joint trajectory as $\mathcal{D}\left(\{{x}_{0:\tau},{u}_{0:\tau}\},\epsilon\right)=\left\{\{{x}^{\prime}_{0:\tau},{u}^{\prime}_{0:\tau}\}\bigg{|}\sum_{t=0}^{\tau}\left(\left\lVert x_{t}-{x}^{\prime}_{t}\right\rVert^{2}_{2}+\left\lVert u_{t}-{u}^{\prime}_{t}\right\rVert^{2}_{2}\right)\leq\epsilon\right\}.$ A local GNE trajectory for $\mathcal{G}$ is then defined as follows.
 Definition 1.
 A set of joint trajectories $\{{x}_{0:\tau}^{i^{\star}},{u}_{0:\tau}^{i^{\star}}\}_{i=1}^{N}$ is a local GNE trajectory for $\mathcal{G}$ if for all agent $i\in\mathcal{N}$, agent i’s trajectory $\{{x}_{0:\tau}^{i^{\star}},{u}_{0:\tau}^{i^{\star}}\}$ is an optimal solution of the following optimization problem withing a local neighborhood $\mathcal{D}\left(\{{x}^{\star}_{0:\tau},{u}^{\star}_{0:\tau}\},\epsilon\right)$ for some $\epsilon>0$:
-$\begin{array}[]{ll}\underset{\{x_{0:\tau}^{i},u_{0:\tau}^{i}\}}{\text{minimize}}&\frac{1}{2}\left\lVert x_{\tau}^{i}-\hat{x}_{\tau}^{i}\right\rVert_{Q^{i}_{\tau}}^{2}+\frac{1}{2}\sum_{t=0}^{\tau-1}\left\lVert x_{t}^{i}-\hat{x}_{t}^{i}\right\rVert_{Q^{i}}^{2}\\ &\qquad\qquad+\frac{1}{2}\sum_{t=0}^{\tau}\left\lVert u_{t}^{i}\right\rVert_{R^{i}}^{2}\\ \mbox{subject to}&x_{0}^{i}=\hat{x}_{0}^{i},\enskip{x}_{t+1}^{i}=f^{i}({x}_{t}^{i},{u}_{t}^{i}),\enskip 0\leq t\leq\tau-1,\\ &g(\tilde{x}_{t},\tilde{u}_{t})\leq{0}_{c},\enskip 0\leq t\leq\tau,\\ \end{array}$ (4)
+$\min_{\{x_{0:\tau}^{i},u_{0:\tau}^{i}\}}$ $\left\lVert x_{\tau}^{i}-\hat{x}_{\tau}^{i}\right\rVert_{Q^{i}_{\tau}}^{2}+\sum_{t=0}^{\tau-1}\left\lVert x_{t}^{i}-\hat{x}_{t}^{i}\right\rVert_{Q^{i}}^{2}+\sum_{t=0}^{\tau}\left\lVert u_{t}^{i}\right\rVert_{R^{i}}^{2}$
+$\mathrm{s.t}\quad$ $x^{i}_{0}=\hat{x}^{i}_{0},\enskip x^{i}_{t+1}=f^{i}(x^{i}_{t},u^{i}_{t}),\enskip 0\leq t\leq\tau-1,$
+$g(\tilde{x}_{t},\tilde{u}_{t})\leq{0}_{c},\enskip 0\leq t\leq\tau,$ (4)
 where we use the following notation in $g(\tilde{x}_{t},\tilde{u}_{t})$
@@
 Intuitively, at a local GNE, no agent can reduce their cost function by independently changing their trajectory to any alternative feasible trajectory within a local neighborhood of the joint equilibrium trajectory, $\{{x}_{0:\tau}^{i^{\star}},{u}_{0:\tau}^{i^{\star}}\}_{i=1}^{N}$.
-Note that in generalized Nash equilibria, the agents’ trajectories are coupled both through their independent objective functions and their joint constraints.
+Note that in CPTGs, the agents’ trajectories are coupled through their joint constraints.
 It is important to highlight that solving (1) requires solving a set of $N$ coupled constrained optimal control problems, which is often computationally expensive.
-Therefore, finding all the local Nash equilibria of the game can be time-consuming when using traditional game theoretic methods that aim to solve coupled optimal control problems.
-Hence, in order to alleviate this problem, we leverage constrained potential games, which we build on to identify different modes of interaction.
-IV Constrained Trajectory Potential Games
-Constrained trajectory potential games are a class of games for which a GNE of the original game can be computed by optimizing a single constrained optimal control problem called Potential, as described in [13].
+Therefore, finding all the local GNEs of the game can be time-consuming when using traditional game theoretic methods that aim to solve coupled optimal control problems.
+However, an important property of CPTGs is that a GNE of the original game can be computed by optimizing a single constrained optimal control problem called Potential, as described in [11].
 This reduces the computational cost of computing the GNEs of the game.
@@
 A set of joint trajectories $\{{x}_{0:\tau}^{i^{\star}},{u}_{0:\tau}^{i^{\star}}\}_{i=1}^{N}$ is a local GNE trajectory if within a local neighborhood $\mathcal{D}\left(\{{x}^{\star}_{0:\tau},{u}^{\star}_{0:\tau}\},\epsilon\right)$, it is an optimal solution of
-$\begin{array}[]{ll}\underset{\{x_{0:\tau},u_{0:\tau}\}}{\text{minimize}}&\frac{1}{2}\left\lVert x_{\tau}-\hat{x}_{\tau}\right\rVert_{Q_{\tau}}^{2}+\frac{1}{2}\sum_{t=0}^{\tau-1}\left\lVert x_{t}-\hat{x}_{t}\right\rVert_{Q}^{2}\\ &\qquad+\frac{1}{2}\sum_{t=0}^{\tau}\left\lVert u_{t}\right\rVert_{R}^{2}\\ \mbox{subject to}&x_{0}=\hat{x}_{0},\enskip x_{t+1}=f(x_{t},u_{t}),\enskip 0\leq t\leq\tau-1,\\ &g(x_{t},u_{t})\leq{0}_{c},\enskip 0\leq t\leq\tau,\end{array}$ (5)
+$\min_{\{x_{0:\tau},u_{0:\tau}\}}$ $\left\lVert x_{\tau}-\hat{x}_{\tau}\right\rVert_{Q_{\tau}}^{2}+\sum_{t=0}^{\tau-1}\left\lVert x_{t}-\hat{x}_{t}\right\rVert_{Q}^{2}+\sum_{t=0}^{\tau}\left\lVert u_{t}\right\rVert_{R}^{2}$
+$\mathrm{s.t}\quad$ $x_{0}=\hat{x}_{0},\enskip x_{t+1}=f(x_{t},u_{t}),\enskip 0\leq t\leq\tau-1,$
+$g(x_{t},u_{t})\leq{0}_{c},\enskip 0\leq t\leq\tau,$ (5)
 for some $\epsilon>0$, where $Q_{\tau}\coloneqq\mathop{\rm blkdiag}(Q^{1}_{\tau},\ldots,Q^{N}_{\tau}),\enskip Q\coloneqq\mathop{\rm blkdiag}(Q^{1},\ldots,Q^{N})$ and $R\coloneqq\mathop{\rm blkdiag}(R^{1},\ldots,R^{N})$.
 Proof.
-This theorem follows from Theorem 2 from [13], which states that if each agent’s cost function depends only on their own states and actions, then the underlying game is a constrained dynamic potential game.
-It is easy to see that agent costs according to (3) depend only on the individual agent’s own states and control inputs.
-Therefore, we can use the results from Theorem 2 from [13] to obtain the potential function of the game $\mathcal{G}$ to be
-$\sum_{i=1}^{N}(\frac{1}{2}\left\lVert x_{\tau}^{i}-\hat{x}_{\tau}^{i}\right\rVert_{Q^{i}_{\tau}}^{2}+\frac{1}{2}\sum_{t=0}^{\tau-1}\left\lVert x_{t}^{i}-\hat{x}_{t}^{i}\right\rVert_{Q^{i}}^{2}+\frac{1}{2}\sum_{t=0}^{\tau}\left\lVert u_{t}^{i}\right\rVert_{R^{i}}^{2}).$ (6)
-Furthermore, using the joint notations, we can re-write the sums in (6) as $\sum_{i=1}^{N}\left(\frac{1}{2}\left\lVert x_{\tau}^{i}-\hat{x}_{\tau}^{i}\right\rVert_{Q^{i}_{\tau}}^{2}\right)=\frac{1}{2}\left\lVert x_{\tau}-\hat{x}_{\tau}\right\rVert_{Q_{\tau}}^{2},$ $\sum_{i=1}^{N}\left(\frac{1}{2}\left\lVert x_{t}^{i}-\hat{x}_{t}^{i}\right\rVert_{Q^{i}}^{2}\right)=\frac{1}{2}\left\lVert x_{t}-\hat{x}_{t}\right\rVert_{Q}^{2},$ $\sum_{i=1}^{N}\left(\frac{1}{2}\left\lVert u_{t}^{i}\right\rVert_{R^{i}}^{2}\right)=\frac{1}{2}\left\lVert u_{t}\right\rVert_{R}^{2}$.
-which in turn gives us (5).
-Therefore, we have proven that $\mathcal{G}$ is a constrained trajectory potential game with potential (5).
-Hence, using Theorem 2 from [13], any local solution of (5) will also be a local GNE of the game $\mathcal{G}$.
+This theorem follows from [11, Thm.
+2].
+Since agent costs from (3) depend only on the individual agent’s own states and control inputs, we use the results from [11, Thm.
+2] to obtain the potential function of $\mathcal{G}$ to be
+$\sum_{i=1}^{N}\left(\left\lVert x_{\tau}^{i}-\hat{x}_{\tau}^{i}\right\rVert_{Q^{i}_{\tau}}^{2}+\sum_{t=0}^{\tau-1}\left\lVert x_{t}^{i}-\hat{x}_{t}^{i}\right\rVert_{Q^{i}}^{2}+\sum_{t=0}^{\tau}\left\lVert u_{t}^{i}\right\rVert_{R^{i}}^{2}\right).$ (6)
+Furthermore, using the joint notations, we can re-write the sums in (6) as (1).
+Therefore, we have proven that $\mathcal{G}$ is a CPTG with potential (1).
+Hence, using Theorem 2 from [11], any local solution of (1) will also be a local GNE of $\mathcal{G}$.
 ∎
-Therefore, using Theorem 1, finding local GNE trajectories is equivalent to finding local solutions to (5).
-The conventional approach to solving the optimization in (5) relies on techniques such as inter-point methods [23] or sequential quadratic programming [24].
+Using Theorem 1, any local solutions of (1) will also be local GNE trajectories.
+We would like to acknowledge that finding local solutions of (1) will not necessarily give us all the GNEs of the original as Theorem 1 provides only a sufficient condition for computing local GNEs.
+Furthermore, it is known that the problem of finding all Nash equilibria is a PPAD-complete problem even in the case of static games [13].
+However, for all practical purposes in multi-agent navigation, finding local solutions of (1) has proven to be enough for us.
+The conventional approach to solving the optimization (1) relies on techniques such as interior point methods [24] or sequential quadratic programming [25].
 Naively, the local solutions can be obtained by exhaustive initialization of random trajectories to an optimization solver, such as IPOPT [24], hoping to explore the solution space of local Nash equilibria.
-However, random exploration using gradient-based methods becomes computationally prohibitive for large $N$ and inefficient as the same local equilibria can be arrived at from multiple random initializations.
-To address this, we utilize a Bayesian inference perspective in the following section to recover all local minima of (5) in an efficient manner.
-V Identifying Local Nash Equilibria via Bayesian Inference
+However, randomly exploring the nonconvex optimization landscape via gradient-based methods is computationally prohibitive and inefficient, as the same local equilibria can be arrived at from multiple random initializations.
+To address this, we utilize a Bayesian inference approach in the following section to efficiently recover all local minima of (1).
+IV Identifying Local Nash Equilibria via Bayesian Inference
 In this section, we propose our approach to solve (1) from a Bayesian inference perspective.
 We reformulate the optimal control problem (1) into an equivalent Bayesian inference problem by utilizing the reference trajectory and constraint violations as virtual measurements that provide evidence for the optimal inference of the system trajectory in (1).
-At its core, the Bayesian inference problem is about approximating the probability density function (PDF) of the trajectories given measurement information derived from the objective and constraints in (5).
-Then, the high probability regions of the PDF to be approximated will correspond to the local GNEs of interest.
-This enables using the sample efficient implicit particle filter [30], which is designed to sample from the high probability regions of the PDF.
-The resulting identified equilibria then serve as a warm-start initialization for a downstream optimizer such as IPOPT, which finds the fine-tuned trajectory.
-We first formulate the MulitNash-PF problem and then outline the algorithm that realizes the framework.
-V-A MultiNash-PF framework via Bayesian Inference
-Given problem (5), we view the objective function and constraints as a source of optimality information to facilitate the inference of the optimal trajectory.
-This motivates the introduction of a state-space model that sets us up to construct the Bayesian reformulation as a joint state and control inference via the objective function and constraint as measurements.
-We construct the virtual system as
-$x_{t+1}$ $=f(x_{t},u_{t}),$ $y_{x,t}$ $=x_{t}+v_{t},$
-$u_{t+1}$ $=w_{t},$ $y_{g,t}$ $=\psi\left(g\left(x_{t},u_{t}\right)\right)+\eta_{t},$ (7)
-for $0\leq t\leq\tau$, where $x_{t}$ and $u_{t}$ are the states of the virtual system, $y_{x,t}$, $y_{g,t}$ are measurements, and $v_{t}$, $w_{t}$, and $\eta_{t}$ are bounded disturbances.
-The function $\psi(\cdot)$ represents a barrier function that can capture the degree of constraint violation by a trajectory.
-We chose the softplus barrier function for its continuity in the forthcoming filtering algorithm.
-The solfplus function is expressed as:
-$\psi(g\left(x_{t},u_{t}\right))=\ln\left(1+\exp(g(x_{t},u_{t}))\right)/\alpha,$ (8)
-where $\alpha$ is a tuneable parameter that controls the strictness of enforcing constraints. Considering the virtual system (V-A), we relax the problem (5) using soft constraints as:
-$\begin{array}[]{ll}\underset{\{v_{t},w_{t},\eta_{t}\}_{t=0}^{\tau}}{\min}&\frac{1}{2}\left\lVert v_{\tau}\right\rVert_{Q_{\tau}}^{2}+\frac{1}{2}\sum_{t=0}^{\tau-1}\left\lVert v_{t}\right\rVert_{Q}^{2}\\ &+\frac{1}{2}\sum_{t=0}^{\tau}\left\lVert w_{t}\right\rVert_{R}^{2}+\frac{1}{2}\sum_{t=0}^{\tau}\left\lVert\eta_{t}\right\rVert^{2}_{Q_{\eta}},\\ \quad\;\mathrm{s.t.}&\begin{bmatrix}x_{0}\\ u_{0}\end{bmatrix}=\begin{bmatrix}\hat{x}_{0}\\ w_{-1}\end{bmatrix},\begin{bmatrix}x_{t+1}\\ u_{t+1}\end{bmatrix}=\begin{bmatrix}f(x_{t},u_{t})\\ 0_{m}\end{bmatrix}+\begin{bmatrix}0_{n}\\ w_{t}\end{bmatrix},\vspace{0.2cm}\\ &\begin{bmatrix}y_{x,t}\\ y_{g,t}\end{bmatrix}=\begin{bmatrix}x_{t}\\ \phi\left(g(x_{t},u_{t})\right)\end{bmatrix}+\begin{bmatrix}v_{t}\\ \eta_{t}\end{bmatrix},0\leq t\leq\tau,\end{array}$ (9)
-where $Q_{\eta}$ is a positive definite weight matrix that assigns the importance of constraint satisfaction.
-Problem (9) differs from (5) in that it considers a soft inequality constraint using the barrier function (8).
-Further, (9) is also known as the moving horizon estimation problem [31] which aims to find the minima of the disturbances $\left\{v_{t}^{\star},w_{t}^{\star},\eta_{t}^{\star}\right\}_{t=0}^{\tau}$.
-A direct attempt at solving (9) using traditional gradient-based methods poses a similar computational burden in identifying local Nash equilibria compared to problem (5).
-However, the above connection between (5) and the moving horizon estimation problem (9) highlights the potential to treat the problem from a Bayesian inference perspective.
-It can be seen that the constraints in (9) are the same as the considered virtual system in (V-A), suggesting that state inference can be performed via efficient estimation methods.
-To proceed, we rewrite (V-A) compactly as
+To this end, we first reformulate problem (1) by considering soft-constraints and introducing auxiliary variables $y_{x,t}$ and $y_{g,t}$ as:
+$\min_{\{v_{t},w_{t},\eta_{t}\}_{t=0}^{\tau}}$ $\left\lVert v_{\tau}\right\rVert_{Q_{\tau}}^{2}+\sum_{t=0}^{\tau-1}\left\lVert v_{t}\right\rVert_{Q}^{2}+\sum_{t=0}^{\tau}\left\lVert w_{t}\right\rVert_{R}^{2}+\left\lVert\eta_{t}\right\rVert^{2}_{Q_{\eta}},$
+$\mathrm{s.t}\quad$ $\begin{bmatrix}x_{t+1}\\ u_{t+1}\end{bmatrix}=\begin{bmatrix}f(x_{t},u_{t})\\ 0_{m}\end{bmatrix}+\begin{bmatrix}0_{n}\\ w_{t}\end{bmatrix},$
+$\begin{bmatrix}y_{x,t}\\ y_{g,t}\end{bmatrix}=\begin{bmatrix}x_{t}\\ \psi\left(g(x_{t},u_{t})\right)\end{bmatrix}+\begin{bmatrix}v_{t}\\ \eta_{t}\end{bmatrix},$
+$\begin{bmatrix}x_{0}\\ u_{0}\end{bmatrix}=\begin{bmatrix}\hat{x}_{0}\\ w_{t-1}\end{bmatrix},\quad 0\leq t\leq\tau,$ (7)
+where $v_{t}$, $w_{t}$, and $\eta_{t}$ are bounded disturbances, and $Q_{\eta}\in\mathbb{R}^{n_{c}\times n_{c}}$ assigns the importance of constraint satisfaction.
+Further, $\psi$ is a barrier function used to quantify the degree of constraint violation by defining
+$\psi\left(g(x_{t},u_{t})\right)=\frac{1}{\alpha}\ln\left(1+\exp\left(g\left(x_{t},u_{t}\right)\right)\right),$ (8)
+where $\alpha$ controls the strictness enforcing constraints.
+We note that the problem in (IV) resembles a moving horizon estimation problem [32].
+A direct attempt at solving (IV) using traditional gradient-based methods poses a similar computational burden in identifying all of the local equilibria compared to problem (1).
+However, the above connection between (1) and the moving horizon estimation problem (IV) highlights the potential to treat the problem from a Bayesian inference perspective.
+To this end, we construct the following virtual state-space model from (IV) as
 $\bar{x}_{t+1}=\bar{f}(\bar{x}_{t})+\bar{w}_{t},\;{\bar{y}}_{t}=\bar{h}(\bar{x}_{t})+\bar{v}_{t},$ (9)
@@
 $\bar{x}_{t}$ $=\begin{bmatrix}x_{t}\\ u_{t}\end{bmatrix},\>\bar{y}_{t}=\begin{bmatrix}y_{x,t}\\ y_{g,t}\end{bmatrix},\>\bar{w}_{t}=\begin{bmatrix}0_{n}\\ w_{t}\end{bmatrix},$
-$\bar{v}_{t}=\begin{bmatrix}v_{t}\\ \eta_{t}\end{bmatrix},\;\bar{f}($ $\bar{x}_{t})=\begin{bmatrix}f(x_{t},u_{t})\\ {0}_{m}\end{bmatrix},\;\bar{h}(\bar{x}_{t})=\begin{bmatrix}x_{t}\\ \psi(g\left(x_{t},u_{t}\right))\end{bmatrix}.$
-Here, we relax $\bar{w}_{t}$ and $\bar{v}_{t}$ for $0\leq t\leq\tau$ to be stochastic disturbances modeled as normally distributed random variables with PDF
-$\bar{v}_{t}\sim\mathcal{N}(0_{n+c},\bar{Q}),\;\bar{v}_{\tau}\sim\mathcal{N}(0_{n+c},\bar{Q}_{\tau}),\;\bar{w}_{t}\sim\mathcal{N}(0_{n+m},\bar{R}),$ (11)
+$\bar{v}_{t}=\begin{bmatrix}v_{t}\\ \eta_{t}\end{bmatrix},\;\bar{f}($ $\bar{x}_{t})=\begin{bmatrix}f(x_{t},u_{t})\\ {0}_{m}\end{bmatrix},\;\bar{h}(\bar{x}_{t})=\begin{bmatrix}x_{t}\\ \psi(g\left(x_{t},u_{t}\right))\end{bmatrix},$
+where we relax $\bar{w}_{t}$ and $\bar{v}_{t}$ for $0\leq t\leq\tau$ to be stochastic disturbances modeled as Gaussian-distributed random variables with density
+$\bar{v}_{t}\sim\mathcal{N}(0_{n+c},\bar{Q}),\;\bar{v}_{\tau}\sim\mathcal{N}(0_{n+c},\bar{Q}_{\tau}),\;\bar{w}_{t}\sim\mathcal{N}(0_{n+m},\bar{R}),$
 with $\bar{Q}\coloneqq\mathop{\rm blkdiag}(Q^{-1},Q_{\eta}^{-1})$, $\bar{Q}_{\tau}\coloneqq\mathop{\rm blkdiag}(Q_{\tau}^{-1},Q_{\eta}^{-1})$, $\bar{R}\coloneqq\mathop{\rm blkdiag}(0_{n\times n},R^{-1})$.
 The state estimation relevant to (9) is to estimate the state $\bar{x}_{t}$ given the measurements $\bar{y}_{t}$.
-We must set the virtual measurement’s value to correspond to the objective of the optimization in (5).
+The value of the virtual measurements must be set in a way that steers the inference problem to have the same optima as in problem (1).
 We achieve this by noting that $y_{x,t}$ is a measurement value used to correct the state estimate of $x_{t}$ towards the reference $\hat{x}_{t}$.
 While $y_{g,t}$ is to satisfy constraints through (8).
-Further, we observe that the barrier function in (8) outputs zero when the constraints are satisfied.
+Observing that the barrier function in (8) outputs zero when the constraints are satisfied, we set the virtual measurement for the barrier function equal to $0_{c}$.
 Hence, the virtual measurement that will drive $\bar{x}_{t}$ towards optimality takes the value
 $\bar{y}_{t}=\begin{bmatrix}\hat{x}_{t}\\ {0}_{c}\end{bmatrix},0\leq t\leq\tau.$
-The Bayesian inference problem pertaining to (10) is to characterize the conditional PDF $p(\bar{x}_{0:\tau}\,|\,\bar{y}_{0:\tau})$.
-In the following, lemma (1) elucidates that characterizing the maximum a posteriori estimate of $p(\bar{x}_{0:\tau}\,|\,\bar{y}_{0:\tau})$ holds equivalence to solving (5).
+The Bayesian inference problem pertaining to (9) is one that characterize the conditional distribution $p(\bar{x}_{0:\tau}\,|\,\bar{y}_{0:\tau})$.
+In the following, lemma (1) elucidates that equivalence holds between a posteriori estimate of $p(\bar{x}_{0:\tau}\,|\,\bar{y}_{0:\tau})$ and the problem in (1).
 Lemma 1.
@@
 $\bar{x}_{0:\tau}^{\star}=\arg\max_{\bar{x}_{0:\tau}}\>\log p(\bar{x}_{0:\tau}\,|\,\bar{y}_{0\tau}).$
-We refer the interested reader to [12, Thm.
+We refer the interested reader to [16, Thm.
 1] for the proof.
@@
 Generally, state estimation of a multi-modal PDF does not admit closed-form solutions [33].
-Hence, we reside to approximate solutions.
+Hence, we resort to approximate solutions.
 A powerful approach is to use Monte Carlo methods that empirically approximate the distribution via a set of $J$ particles as
@@
 where $\delta(\cdot)$ is the dirac delta function and $w_{\tau}^{i}$ is the weight assigned to a sample trajectory $\bar{x}_{0:\tau}^{i}$ for $j=1,\ldots,J$.
-After identifying the local GNE through the above approximation, we extract the trajectories that appear as local GNE and use them to initialize an optimization solver.
-This completes the exposition of the MultiNash-PF framework.
-We employ the implicit particle filter  [30] due to its two-fold benefit for implementing the MultiNash-PF framework.
-First, as suggested in lemma 1, the local maximums (i.e high-probability regions) of $p(\bar{x}_{0:\tau}\,|\,\bar{y}_{0:\tau})$ correspond to the local GNEs.
-Hence, focusing our sampling efforts towards these high-probability regions can reduce the required samples for the approximation in (12).
+To achieve the approximation in (10), we employ the implicit particle filter  [31] due to the twofold benefit it holds to implement the MultiNash-PF framework.
+First, as suggested in lemma 1, the local maximums (i.e., high-probability regions) of $p(\bar{x}_{0:\tau}\,|\,\bar{y}_{0:\tau})$ correspond to the local GNEs.
+Hence, focusing our sampling efforts towards these high-probability regions can reduce the number of required samples for the approximation in (10).
+This approach aligns precisely with the principle of implicit importance sampling, whose objective is to build a mapping hat draws particles from the high-probability regions of the filtering distribution in (10).
 Second, we want to avoid the traditionally needed resampling step in particle filtering methods to ensure that the inferred trajectories are smooth.
-Fortunately, the implicit particle filter satisfies both of the aforementioned criteria.
-We leverage the Markovian property of the virtual system (10) to approximate (12).
-Using the Bayes’ rule, we have
+Fortunately, the implicit particle filter does not suffer from the problem of particle degeneracy, and filtering can be achieved by disregarding resampling.
+We leverage the Markovian property of the virtual system (9) to recursively approximate (10). Using Bayes’ rule, we have
 $p(\bar{x}_{0:t}\,|\,\bar{y}_{0:t})=p(\bar{y}_{t}\,|\,\bar{x}_{t})p(\bar{x}_{t}\,|\,\bar{x}_{t-1})p(\bar{x}_{0:t-1}\,|\,\bar{y}_{0:t-1}),$ (11)
 for $0\leq t\leq\tau$.
-This indicates that given the particle $\bar{x}_{t-1}^{j}$ from the prior distribution $p(\bar{x}_{0:t-1}\,|\,\bar{y}_{0:t-1})$, we can recursively obtain $\bar{x}_{t}^{j}$, reducing the complexity involved in directly sampling from a high-dimensional PDF.
-As shown in [30], we use implicit importance sampling and identify an implicit mapping to sample from high probability regions of (13).
-For this purpose, we make a local Gaussian assumption around $\bar{x}_{t-1}^{j}$ such that
-$p(\bar{x}_{t}^{j}|\bar{y}_{0:t})\sim\mathcal{N}\left(\hat{m}_{t}^{j},\hat{\Sigma}_{t}^{x,j}\right),$ (14)
-where $\hat{m}_{t}^{j}$ and $\hat{\Sigma}_{t}^{x,j}$ are the mean and error covariance of $\bar{x}_{t}^{j}$, respectively.
-These quantities are then found recursively by using a nonlinear Kalman filter.
-Then, the implicit map $X(\gamma_{t}^{j})$ can be computed in closed form as
-$\bar{x}_{t}^{j}=\hat{m}_{t}^{j}+\sqrt{\hat{\Sigma}_{t}^{x,j}}\gamma_{t}^{j},$ (15)
-where $\gamma_{t}^{j}$ is a sample from $\mathcal{N}(0,I_{n})$ and $I_{n}$ denotes the identity matrix of size $n$.
+This indicates that given the particle $\bar{x}_{t-1}^{j}$ from the prior distribution $p(\bar{x}_{0:t-1}\,|\,\bar{y}_{0:t-1})$, we can obtain $\bar{x}_{t}^{j}$.
+As shown in [31], we use implicit importamapping sampling and identify an implicit mapping to sample from high probability regions of (11).
+The implicit map is constructed by considering sampling from a reference random vector $\xi_{t}$ with density $p(\gamma_{t})$ and map it to high probability regions of $p(x_{0:\tau}\,|\,y_{0:\tau})$, such that a sample $\gamma^{i}_{t}$ is mapped to $x_{t}^{i}$.
+The mapping connects the highly probable regions of $p(\xi)$ to $p(x_{0:\tau}\,|\,y_{0:\tau})$ by letting
+$F(x_{t}^{i})-\min F(x_{t}^{i})=G(\gamma_{t}^{i})-\min G(\gamma_{t}^{i}),$ (12)
+where $F(x_{t}^{i})=-\log\left(p(x_{t}^{i}\,|\,y_{1:t})\right)$ and $G(\gamma_{t}^{i})=-\log\left(p(\gamma_{t}^{i})\right)$.
+Finding an explicit solution to (12) is computationally intractable due to the involved nonconvex optimization.
+To bypass this computational challenge, a local Gaussian approximation around $\bar{x}_{t-1}^{j}$ is employed such that
+$p(\bar{x}_{t}^{j}|\bar{y}_{0:t})\sim\mathcal{N}\left(\hat{m}_{t}^{j},\hat{\Sigma}_{t}^{x,j}\right),$
+where $\hat{m}_{t}^{j}$ and $\hat{\Sigma}_{t}^{x,j}$ are the mean and covariance of $\bar{x}_{t}^{j}$, respectively.
+These quantities can be approximated recursively using an Unscented Kalman filter (UKF) [34] for all $j=1,\ldots,N$.
+Then, the implicit map can be computed in closed form as
+$\bar{x}_{t}^{j}=\hat{m}_{t}^{j}+\sqrt{\hat{\Sigma}_{t}^{x,j}}\gamma_{t}^{j},$
+where $\gamma_{t}^{j}$ is a sample from $p(\gamma_{t})\sim\mathcal{N}(0,I_{n})$ and $I_{n}$ denotes the identity matrix of size $n$.
+This process is repeated until $t=\tau$ for each particle to obtain the set of trajectories $\{\bar{x}_{0:\tau}\}_{j=1}^{J}$, which will include the coarse estimates of local GNEs.
+This completes the unscented implicit particle filter, identifying the coarse estimates of multiple local Nash equilibria in (1).
+In the last step, the local equilibria are extracted to be from $\{x_{0:\tau}\}_{j=1}^{J}$.
+As viewed in Fig.
+4, the particle set contains multiple trajectories for each GNE.
+Therefore, to isolate the GNE trajectories, a clustering method is used to aggregate the particles close to each other in one cluster.
+One can leverage methods such as hierarchical clustering [35] or DBSCAN [36].
+It should be noted that clustering methods require a metric to compute distances between the data points.
+The Fréchet distance [37] is utilized as our metric to compute distances between the trajectories.
+The clustering of trajectories will result in one cluster for each GNE by taking the average of trajectories over each cluster and feed them as a warm start to the IPOPT solver to obtain the local GNEs.
+The algorithmic steps of the proposed MultiNash-PF are provided in Algorithm 1.
 Algorithm 1 MultiNash-PF: Multiple local GNEs based on particle filter
@@
 4:Initialize the particles: $\bar{x}_{0}^{j}$, $\hat{\Sigma}^{x,j}_{0}$, and $w_{0}^{j}$ for $j=1,\ldots,J$
-5:for $t=0,1,\ldots,\tau-1$ do
-6:     for $i=1,2,\ldots,J$ do
-7:Run Kalman prediction step:
-8:         $\left(m_{t+1}^{j},\Sigma^{x,j}_{t+1}\right)\leftarrow\text{{\bf UT}}(\bar{x}_{t}^{j},\Sigma_{t}^{x,j},\bar{R},\bar{f})$
-9:Run Kalman update step:
-10:         if $t\leq\tau-2$ then
-11:              $(\hat{\bar{y}}_{t+1}^{j},\Sigma^{y,j}_{t+1},\Sigma^{xy,j}_{t+1})\leftarrow\text{{\bf UT}}(m_{t+1}^{j},\Sigma_{t+1}^{x,j},\bar{Q},\bar{h})$
-12:         else
-13:              $(\hat{\bar{y}}_{t+1}^{j},\Sigma^{y,j}_{t+1},\Sigma^{xy,j}_{t+1})\leftarrow\text{{\bf UT}}(m_{t+1}^{j},\Sigma_{t+1}^{x,j},\bar{Q}_{\tau},\bar{h})$
-14:         end if
-15:         $K_{t+1}^{j}\leftarrow\Sigma_{t+1}^{xy,j}(\Sigma_{t+1}^{y,j})^{-1}$
-16:         $\hat{m}^{x,j}_{t+1}=m_{t+1}^{j}+K_{t+1}^{j}\left(\bar{y}_{t+1}-\hat{\bar{y}}_{t+1}^{j}\right)$
-17:         $\hat{\Sigma}_{t+1}^{x,j}\leftarrow\Sigma_{t+1}^{x,j}-K_{t+1}^{j}\Sigma^{y,j}_{t+1}(K_{t+1}^{j})^{\top}$
-18:Draw a high-probability sample $\gamma_{t+1}^{j}\sim\mathcal{N}(0,I_{n})$ for the implicit mapping in (15):
-19:         $\bar{x}_{t+1}^{j}\leftarrow\hat{m}_{t+1}^{x,j}+\sqrt{\hat{\Sigma}_{t+1}^{x,j}}\gamma_{t+1}^{j}$
-20:Weight and normalize samples:
-21:         $w_{t+1}^{j}=\frac{w_{t}^{j}p(\bar{y}_{t+1}\,|\,\bar{x}_{t}^{j})}{\sum_{j=1}^{J}w_{t}^{j}p(\bar{y}_{t+1}\,|\,\bar{x}_{t}^{j})},\qquad p(\bar{y}_{t+1}\,|\,\bar{x}_{t}^{j})\sim\mathcal{N}(\hat{\bar{y}}_{t+1}^{j},\Sigma^{y,j}_{t+1})$
-22:     end for
-23:     Perform resampling if necessary
-24:end for
-25:Extract trajectories from $\{x_{0:\tau}\}_{j=1}^{J}$ and warm start IPOPT solver to obtain the local GNEs $\{{x}_{0:\tau}^{i^{\star}},{u}_{0:\tau}^{i^{\star}}\}_{i=1}^{N}$ of $\mathcal{G}$.
-V-B Implicit Particle Filter via Unscented Kalman filter
-As in [30], we note that the local Gaussian approximation in (14) can be recursively computed through a nonlinear Kalman filter evolving according to (13) for all $j=1,\ldots,J$.
-Our nonlinear Kalman filter of choice is the unscented Kalman filter due to its high filtering accuracy for low-dimensional state space models [33].
-The unscented Kalman filter is based on the unscented transform, which utilizes deterministic samples, known as sigma points, to trace the evolution of the mean and covariance of a Gaussian random variable through a nonlinear function.
-The unscented Kalman filter [33] utilizes the unscented transform in the Kalman prediction and measurement update steps to keep track of the Gaussian statistics through the state transition and measurement function in (10).
-From now on, we use UT to denote Unscented Transform algorithm.
-Using the unscented Kalman filter, we can recursively compute the local Gaussian approximation (14) at time $t$.
-Then, we implicitly map the particle to the high-probability regions by sampling $\gamma_{t}^{i}\sim\mathcal{N}(0,I_{n})$ for (15) to obtain the particle $\bar{x}_{t}^{i}$.
-If necessary, we perform resampling for the particle trajectory $\bar{x}_{0:t}^{i}$.
-This process is repeated until $t=\tau$ for each particle to obtain the set of trajectories $\{\bar{x}_{0:\tau}\}_{j=1}^{J}$, which will include the coarse estimates of local GNEs.
-This completes the unscented implicit particle filter, which has identified the coarse estimates of multiple local Nash equilibria in (5).
-As the last step, the estimates of local equilibria are extracted and used as warm start initialization to the IPOPT optimizer to find the optimal local GNEs.
-To extract the equilibria, we carefully observe the trajectories and select the trajectories that are repeatedly taken by multiple particles.
-The pragmatic extraction of equilibria from the particle trajectories is nontrivial and is left for future study.
-The implementation details of the proposed MultiNash-PF through the unscented implicit particle filter are provided in Algorithm 1.
-VI Numerical Simulations
+5:Employ Algorithm-1 from [31] combined with UKF to obtain coarse solutions, $\{x_{0:\tau}\}_{j=1}^{J}$.
+6:Extract trajectories through clustering from $\{x_{0:\tau}\}_{j=1}^{J}$ and warm start IPOPT solver to obtain the local GNEs $\{{x}_{0:\tau}^{i^{\star}},{u}_{0:\tau}^{i^{\star}}\}_{i=1}^{N}$ of $\mathcal{G}$.
+V Numerical Simulations
 To showcase the capabilities of our method, we take motivation from real-life highway scenarios.
@@
 where $d(x^{1}_{t},x^{2}_{t})\coloneq\sqrt{(p^{1}_{t}-p^{2}_{t})^{2}+(q^{1}_{t}-q^{2}_{t})^{2}}$ is the Euclidean distance between the two agents and $r_{\text{col}}\coloneq 3m$ is the collision avoidance radius between the two agents.
-With this setup, we employ MultiNash-PF from Algorithm 1 to efficiently capture the multi-modality in the solution space and then extract the solutions to provide a warm start trajectory to a solver like IPOPT [23] to recover different local GNEs.
-We use the Julia programming language and utilize the JuMP [34] library to solve the game using IPOPT solver.
+With this setup, we employ MultiNash-PF from Algorithm 1 to efficiently capture the multi-modality in the solution space and then extract the solutions using hierarchical clustering [35] to provide a warm start trajectory to IPOPT [24] solver to recover different local GNEs.
+We use the Julia programming language and utilize the JuMP [38] library to solve the game using the IPOPT solver.
 The results of our method are provided in Fig 2.
-As we can see, the presence of two agents and collision avoidance constraints induces two different possible Nash equilibria.
+As we can see, the presence of two agents and collision avoidance constraints induces two different possible GNEs.
 Either both the agents will yield to their right, or they will yield to their left to avoid collision.
@@
 It is challenging to recover all six equilibria without utilizing the capabilities of the implicit particle filter.
-Upon performing simulation experiments with our method, we observe that MultiNash-PF is able to recover all six Nash equilibria as shown in Fig.
-2.
+Upon performing simulation experiments with our method, we observe that MultiNash-PF can recover all six Nash equilibria as shown in Fig.
+4 and Fig.
+3.
 (a) Histogram of the solve time comparison of MultiNash-PF and the baseline.
 (b) Histogram of the number of the IPOPT runs required to obtain all six equilibria.
-Figure 3: Histograms of the comparison of MultiNash-PF with the baseline for the experiment shown in Fig. 2.
+Figure 5: Comparison of MultiNash-PF with the baseline for the experiment shown in Fig. 3.
 (a) The average solve time for MultiNash-PF is $3.36\pm 0.59s$ while the baseline’s average solve time is $6.71\pm 2.83s$.
@@
 MultiNash-PF requires 3 times fewer IPOPT runs.
-VI-1 Comparison with Baseline
+V-1 Comparison with Baseline
 In order to compare the effectiveness of our method, we consider a baseline method.
 In the baseline method, we directly initialize the IPOPT solver with random perturbations of reference trajectories and compute the resulting optimal solutions.
-For our method, we fixed $J=50$ particles and obtained the solution trajectories using implicit particle filtering from Algorithm 1 described in Section V.
+For our method, we fixed $J=50$ particles and obtained the solution trajectories using implicit particle filtering from Algorithm 1 described in Section IV.
 After that, we extract all the different equilibria and provide them as a warm starting point for IPOPT to obtain the final solutions.
@@
 A histogram of the number of IPOPT runs required for obtaining all six equilibria is plotted in Fig.
-3(b).
+5(b).
 Furthermore, the solve time to recover all 6 different Nash equilibria for the baseline is $6.71\pm 2.83s$, while for MultiNash-PF, the solve time is $3.36\pm 0.59s$.
@@
 A histogram of the solve times of the baseline compared to our method is plotted in Fig.
-3(a).
+5(a).
 As explained earlier, our method’s solve time is a combination of the time used in implicit particle filtering and the time used by IPOPT.
@@
 Therefore, in our MultiNash-PF, the time used by implicit particle filtering is $11\%$ of the total solve time, which shows that filtering is very efficient in exploring the solution space of local Nash equilibria and contributes really less to the overall computation time of MultiNash-PF.
+VI Hardware Experiment
+Finally, we conduct a hardware experiment where a human and a robot exchange their position as shown in Fig.
+1.
+We perform these experiments using a TurtleBot 4 which is a ground robot with unicycle dynamics.
+We assume that the human also follows unicycle dynamics.
+The collision avoidance constraint between the TurtleBot and human is assumed to be the same as of (14).
+We use the Vicon motion capture system to measure the current position of both the robot and the human.
+Similar as shown in Fig.
+2, two GNEs may exist which will correspond to two different modes of interactions.
+Namely, either the human and robot yield to their right or they yield to their left.
+For the robot to safely navigate around the human without collision, the robot needs to be aware of both the GNEs and needs to be able to determine which GNE the human chooses.
+Using MultiNashPF, the robot first uses an implicit particle filter to obtain the coarse estimates of the two different equilibria and then utilizes them as a warm start for IPOPT to obtain two modes of interaction.
+We assume that the interaction happens over $T=10s$.
+Then similar to [11], the robot uses a potential function based game theoretic planner in a model-predictive fashion with a planning horizon of 5s and $\Delta t=0.1s$ to plan its motion.
+The robot faces two choices of GNEs to follow recovered from the MultiNashPF.
+At any given time $t$ of planning, the robot computes the Fréchet distance [37] between the trajectory exhibited by the human so far and the human’s trajectory in both the computed GNEs.
+Let the distances of human trajectory from both the GNEs be denoted by $d_{1}$ and $d_{2}$, respectively.
+The robot keeps an uncertainty on the GNE that the human chooses until $|d_{1}-d_{2}|>d^{*}$ where $d^{*}$ is the specified threshold.
+For example, in Fig.
+1 at $t=2$, values of $d_{1}$ and $d_{2}$ are almost identical which indicates that human has not chosen their mode of interaction.
+Whenever $|d_{1}-d_{2}|>d^{*}$ occurs, the robot becomes certain of the mode of interaction followed by the human.
+Then the robot follows GNE corresponding to $\arg\min_{i\in\{1,2\}}\{d_{i}\}$.
+For example, in Fig.
+1, at $t=4$, the Fréchet distance of the human from its first GNE trajectory is very high while for the second GNE, the Fréchet distance is low indicating that the human has chosen the second GNE which corresponds to both the agents yielding to their right.
+Using our method, the robot autonomously identifies this mode, moves accordingly, and reaches its goal location while successfully avoiding collision with the human.
+We ran 10 different trials and each time we asked the human to either yield to their right or left randomly.
+Our algorithm was able to successfully compute both modes and identify the mode that the human was following.
+Using our algorithm, the robot was able to successfully avoid collision with the human in real-time and reach its goal location in all 10 runs.
 VII Conclusion
-In this work, we introduced MultiNash-PF, a novel algorithm designed to efficiently compute multiple local generalized Nash equilibria (GNEs) in constrained trajectory games.
-By integrating potential game theory with implicit particle filtering, MultiNash-PF provides coarse estimates of local minimizers, which are then refined using off-the-shelf optimization solvers.
-Our experimental results in two-player trajectory games involving nonlinear dynamics and collision avoidance constraints demonstrate that MultiNash-PF not only identifies all local Nash equilibria but also significantly reduces computation time compared to traditional random initialization methods.
-This enhanced efficiency and accuracy underline the potential of MultiNash-PF to improve identifying various interaction modes in complex, real-world environments, paving the way for more robust and socially acceptable autonomous systems.
+In this work, we introduced MultiNash-PF, a novel algorithm for efficiently computing multiple local GNEs in CPTGs.
+By leveraging potential game theory and implicit particle filtering, MultiNash-PF efficiently identifies coarse estimates of local minimizers, which are then refined using optimization solvers to obtain distinct interaction modes.
+Our numerical simulations demonstrate that MultiNash-PF significantly reduces computation time—by up to 50% compared to baseline methods—while effectively capturing the multimodal nature of multi-agent interactions.
+Furthermore, our real-world human-robot interaction experiments highlight the algorithm’s ability to reason about multiple interaction modes and resolve conflicts in real-time.
+These results underscore the potential of MultiNash-PF to enhance interactive motion planning, enabling more adaptive and socially aware autonomous systems.

### v3

@@
 Abstract
-Modern robotic systems frequently engage in complex multi-agent interactions, many of which are inherently multi-modal, meaning they can lead to multiple distinct outcomes.
+Modern robotic systems frequently engage in complex multi-agent interactions, many of which are inherently multi-modal, i.e., they can lead to multiple distinct outcomes.
 To interact effectively, robots must recognize the possible interaction modes and adapt to the one preferred by other agents.
-In this work, we propose an efficient algorithm for capturing the multimodality in multi-agent interactions.
-We leverage a game-theoretic planner to model interaction outcomes as equilibria where each equilibrium corresponds to a distinct interaction mode.
-We then develop an efficient algorithm to identify all the equilibria, allowing robots to reason about multiple interaction modes.
-More specifically, we formulate interactive planning as Constrained Potential Trajectory Games (CPTGs) and model interaction outcomes by local Generalized Nash equilibria (GNEs) of the game.
-CPTGs are a class of games for which a local GNE can be found by solving a single constrained optimal control problem where a potential function is minimized.
+In this work, we propose MultiNash-PF, an efficient algorithm for capturing the multimodality in multi-agent interactions.
+We model interaction outcomes as equilibria of a game-theoretic planner, where each equilibrium corresponds to a distinct interaction mode.
+Our framework formulates interactive planning as Constrained Potential Trajectory Games (CPTGs), in which local Generalized Nash Equilibria (GNEs) represent plausible interaction outcomes.
 We propose to integrate the potential game approach with implicit particle filtering, a sample-efficient method for non-convex trajectory optimization.
@@
 I Introduction
-Modern-world robotics applications are inherently multi-agent in nature.
-For instance, autonomous cars on highways need to account for other vehicles while planning their motion.
-Most of the real-world interactions are multi-modal.
+Many real-world robotic interactions are inherently multi-modal.
 For example, in the two-player interaction illustrated in Fig.
@@
 Here, two outcomes exist: both agents yielding to their right or left.
-If the two agents were to act independently, they could reach different outcomes, which would result in a collision.
+If the two agents act independently, they might choose conflicting actions.
+For example, one yields to the left and the other to the right, which can potentially result in a collision.
 What trajectory each agent would prefer to pick depends on the convention they might be following.
-For example, in some cultures, in such settings, the convention is to yield to the right to avoid collisions while in other cultures, the convention is to yield to the left.
-Such conventions usually arise out of recurring coordination problems [1].
-Therefore, autonomous agents need to be aware of the existence of multi-modality in the interactions for efficient coordination with people who follow different conventions.
-Some recent works have shown that computing such multiple trajectories is critical to aligning agents’ preferences over different outcomes [2] or helping coordinate agents’ actions via recommendation [3].
+For example, in some cultures, in such settings, the convention is to yield to the right to avoid collisions, while in other cultures, the convention is to yield to the left.
+Such conventions typically emerge from repeated coordination problems [1], and autonomous agents must recognize them and account for the multi-modality of interactions to coordinate with humans effectively.
+Some recent works have shown that reasoning about the multi-modality of interactions is critical to aligning agents’ preferences over different outcomes [2] or helping coordinate agents’ actions via recommendation [3].
+In this paper, we address this challenge by proposing MultiNash-PF, an algorithm that efficiently identifies multiple likely interaction outcomes in multi-agent domains and enables agents to adapt their plans accordingly.
 Figure 1: Snapshot of an interaction between a robot and a human when they exchange positions.
-Using MultiNash-PF the robot can compute both modes of interaction.
+Using MultiNash-PF, the robot can compute both modes of interaction.
 At $t=2$, the robot is uncertain about the preferred mode of the human since the human trajectory’s distance from both equilibria is almost the same.
-At $t=4$, the human trajectory is closer to the second GNE therefore, the robot chooses to follow the second GNE accordingly and avoid collision with the human.
-To formalize this problem, we will use dynamic game theory [4] which has shown to be powerful in developing interactive and socially complaints robots [5, 6, 7, 8].
-We formulate a multi-agent interactive trajectory planning problem as a Constrained Potential Trajectory Game (CPTG).
-In CPTGs, the interaction outcomes are captured through generalized Nash equilibrium (GNE), which are constrained Nash equilibria.
-When multiple agents- such as robots and humans- interact, a GNE corresponds to a joint action such that no agent has incentives to deviate from its equilibrium actions unilaterally.
-Several recent works have shown great success in utilizing GNE for multi-agent motion planning in interactive domains [9, 10, 11, 12].
-A key challenge in interactive domains, however, is the existence of multiple equilibria, many of which are equally valid.
+At $t=4$, the human trajectory is closer to the yield-right GNE; therefore, the robot chooses the corresponding GNE and avoids collision with the human.
+We formulate the problem of computing multi-agent interaction outcomes as a multi-agent trajectory planning problem.
+To formalize interactions, we use dynamic game theory [4], which has been shown to be powerful in developing interactive and socially compliant robots [5, 6, 7, 8] where it has been shown that the interaction outcomes are captured through generalized Nash equilibrium (GNE).
+When multiple agents interact, a GNE corresponds to a joint action such that no agent has incentives to deviate from its equilibrium actions unilaterally.
+Several recent works have shown great success in utilizing GNE for modeling interaction outcomes in interactive domains [9, 10, 11, 12].
+However, there typically exist multiple equilibria in interaction games, many of which are equally valid.
 We argue that each equilibrium corresponds to a distinct mode of interaction.
@@
 Consequently, we propose that in order to capture different interaction modes, robots must reason about multiple interaction equilibria and select the one that aligns with the convention adopted by other agents.
-Computing different local GNEs requires an efficient way to solve games.
-In practice, computing even a single GNE is often challenging [13] because solving for GNE amounts to solving a set of coupled constrained optimal control problems.
-Typically, such a solution requires solving first-order KKT conditions of the underlying problems.
-Therefore, solving the GNE often reduces to addressing a set of complementarity conditions, a process known to be computationally demanding, especially for large-scale applications [14].
-Furthermore, computing all the different local GNEs requires initializing the game solvers with many different initial guesses of the equilibrium trajectories from prior knowledge of the specific scenario.
-This further increases the required number of initial guesses for the game solvers to find all local GNEs.
+However, this is a non-trivial task as it requires the robot to find all local GNEs.
 To our best knowledge, there is no principled approach to compute multiple local Nash equilibria in constrained trajectory games other than exhaustive random initializations of the game solvers.
-Figure 2: The Nash equilibrium trajectories found by MultiNash-PF when two unicycle agents change their positions.
-First, we use the implicit particle filter to discover the coarse estimates of two different equilibria, and then utilize them as a warm starting for IPOPT to obtain the two Nash equilibrium trajectories.
-In this paper, we tackle this challenge by introducing MultiNash-PF, a novel framework designed to efficiently compute multiple local generalized Nash equilibria (GNEs) in constrained trajectory games.
+In this paper, we tackle this challenge by introducing MultiNash-PF, a novel framework designed to efficiently compute multiple local GNEs in constrained trajectory games.
 This, in turn, enables robots to recognize and reason about multiple interaction modes.
-Our approach integrates constrained potential games [15], a class of games that allow for the computation of local GNEs by solving a single constrained optimal control problem, with constraint-aware implicit particle filtering—a sample-efficient method recently applied to nonconvex trajectory optimization [16] to achieve efficient identification of multiple equilibria.
-Using CPTGs, we can compute a local GNE by solving a single constrained optimal control problem using off-the-shelf optimization solvers.
+Our approach integrates 1) constrained potential games [13], a class of games for which local GNEs can be computed by solving a single constrained optimal control problem, and 2) constraint-aware implicit particle filtering which is a sample-efficient method recently applied to nonconvex trajectory optimization [14], to develop a principled approach for identifying various interaction outcomes.
+Using CPTGs, we can compute a local GNE by minimizing the potential function of the game, which is equivalent to solving a single constrained optimal control problem.
+We can solve the resulting optimal control problem using off-the-shelf optimization solvers.
 Then, we use a constraint-aware implicit particle filtering method to identify multiple local minima of the potential minimization problem.
-This method eliminates the requirements of random initialization of the optimization solver by efficiently obtaining coarse estimates of different local GNEs.
+This method eliminates the need for random initialization of the optimization solver by efficiently obtaining coarse estimates of different local GNEs.
 Next, we use the identified minima as initialization for an optimization solver to refine our solution.
-The solver then computes all the different local GNEs of the original game.
-Leveraging the sample efficiency of the implicit particle filtering method, MultiNash-PF can explore and identify the multiple local GNEs faster than the random initialization of optimization solvers.
-We further show the real-time application of MultiNashPF in a human-robot navigation setting where the robot can reason about the existence of different local GNEs and plan its motion autonomously based on the modality the human prefers.
-Figure 3: MultiNash-PF identifies all the local GNE trajectories for a challenging two agent trajectory game where two agents swapped positions while avoiding an obstacle.
-In this scenario, the obstacle radius is greater than the collision avoidance radius of the two agents, leading to the discovery of 6 distinct Nash equilibria.
-MultiNash-PF effectively obtains coarse estimates of local GNEs through implicit particle filtering, which are then refined using IPOPT to obtain local GNEs.
-Figure 4: When using MultiNash-PF on a difficult two agent trajectory game, we first employ an implicit particle filter to obtain coarse estimates of GNEs.
+The solver then computes all the different local GNEs of the original game, which correspond to multiple modes of interaction among the agents.
+Leveraging the sample efficiency of the implicit particle filtering method, MultiNash-PF can explore and identify multiple local GNEs faster than random initialization of optimization solvers.
+We further show the real-time application of MultiNash-PF in a human-robot navigation setting where the robot can reason about the existence of different local GNEs and plan its motion autonomously based on the modality the human prefers.
 II Related Works
 Game Theoretic Planning.
-One of the earliest works laying the foundation of game theoretic motion planning was [17].
-Game theoretic motion planners that seek Stackelberg equilibria were proposed in [18, 7].
-Iterative best response methods for computing Nash equilibria were developed in [19, 20].
-Iterative linear quadratic approximations to compute Nash equilibria were introduced in [6].
-ALGames was developed in [21] which is an augmented lagrangian-based solver for constrained dynamic games.
-A sequential linear-quadratic technique game-based technique to compute GNE was proposed in [22].
-However, these methods are computationally expensive for real-time implementations as the number of agents increases.
-Potential games-based methods were developed later in [15, 23, 11, 12] to reduce the complexity of computing equilibria.
+Game-theoretic planning has been widely used for modeling interactive domains.
+One of the foundational works in game theoretic motion planning was [15].
+A wide range of methods have been proposed to compute different types of equilibria such as Stackelberg equilibria [16, 7], Nash equilibria, and generalized Nash equilibria through techniques such as best response dynamics [17, 18], terative linear-quadratic approximations [6, 10, 19], augmented Lagrangian-based solvers [20], and potential games-based methods  [13, 21, 11, 12, 22, 23].
 However, the shortcomings of all these methods lie in the fact that they find only one Nash equilibrium of the underlying game and ignore the multi-modal nature of the problem that is crucial for coordination among agents.
 In [2], the authors recognize the existence of multiple equilibria and focus on inferring the equilibrium of other agents in the environment.
-However, the methodology relies on computing the different approximate local Nash by rollouts of randomly selected initial strategies, which we show in this work performs poorly compared to our method.
-Multiple Local Solutions to Constrained Optimal Control.
-In general, optimal control problems are often formulated as nonconvex constrained optimization problems, which are solved through the conventional gradient-based approaches such as inter-point methods [24] and sequential quadratic programming [25].
-However, given an initialization, they can only find a single local optimal solution.
-However, in the present context of identifying all potential local GNEs, the gradient-based method would require numerous initializations to find all the local solutions.
-Departing from gradient-based methods, a Bayesian inference approach has been proposed to solve model predictive control problem [26, 27, 28].
-The approach formulates the optimal control problem from a Bayesian inference perspective and employs particle filtering as its estimation method.
-At its core, particle filtering utilizes sequential Monte Carlo sampling to perform state estimation for highly nonlinear systems [29, 30].
-The efficiency and accuracy of the estimation largely depend on the quality of the sampled particles.
-Recently, in [31], an efficient particle filtering method was proposed based on the principle of implicit sampling.
+However, their methodology relies on computing different approximate local Nash equilibria using rollouts of randomly selected initial strategies, which we show in this work performs poorly compared to our method.
+Multiple Local Solutions to a Constrained Optimal Control Problem.
+Optimal control problems that are relevant to robotics are often nonconvex constrained optimization problems, which are solved using gradient-based methods such as interior-point algorithms [24] and sequential quadratic programming [25].
+These methods, however, are limited to finding a single local solution per initialization.
+In the context of identifying all local GNEs, this limitation results in a significant computational burden as numerous initial guesses are required to explore the solution space.
+To address this, we depart from traditional approaches and adopt a Bayesian inference framework that enables simultaneous recovery of multiple local GNEs.
+Recent work [26, 27, 28] has demonstrated that constrained optimization problems can be reformulated as Bayesian inference tasks, enabling estimation algorithms to overcome the limitations of gradient-based solvers.
+In this setting, particle filtering methods [29, 30] provide a principled approach to capturing multimodal distributions.
+We leverage this by employing the implicit particle filter, which improves sample efficiency by focusing particles in high-probability regions of the posterior [31].
 III Constrained Potential Trajectory Games
-We formulate the multi-agent interactive trajectory planning problem as an $N$-agent constrained potential trajectory game (CPTG).
 Let $\mathcal{N}:=\{1,2,\ldots,N\}$ denote the set of agents’ indices.
 We assume that the game is played over a finite time horizon of $\tau\in\mathbb{N}$ time-steps, where $\mathbb{N}$ is the set of natural numbers.
-Let $x^{i}_{t}\in\mathbb{R}^{n_{i}}$ and $u_{t}^{i}\in\mathbb{R}^{m_{i}}$ denote the state and control of the $i$-th agent at time $t$, respectively.
+Let $x^{i}_{t}\in\mathbb{R}^{n_{i}}$ and $u_{t}^{i}\in\mathbb{R}^{m_{i}}$ denote the state and control input of the $i$-th agent at time $t$, respectively.
 We assume that the state of the $i$-th agent evolves as follows:
@@
 Therefore, $f:\mathbb{R}^{n}\times\mathbb{R}^{m}\to\mathbb{R}^{n}$ will be a vector-valued dynamics function for the joint system.
-In CPTGs, each agent has to satisfy some constraints that couple different agents’ states and controls.
-Let $g:\mathbb{R}^{n}\times\mathbb{R}^{m}\to\mathbb{R}^{c}$ with $c\in\mathbb{N}$ denote a continuously differentiable function that defines the constraints on the joint state $x_{t}$ and joint control $u_{t}$ at each time:
+We assume that each agent has to satisfy some constraints that couple different agents’ states and control inputs.
+Let $g:\mathbb{R}^{n}\times\mathbb{R}^{m}\to\mathbb{R}^{c}$ with $c\in\mathbb{N}$ denote a continuously differentiable function that defines the constraints on the joint state $x_{t}$ and joint control input $u_{t}$ at each time:
 $g(x_{t},u_{t})\leq{0}_{c},$ (2)
 where ${0}_{c}$ is a vector of zeros in $\mathbb{R}^{c}$.
-The aim of the constraint function $g$ is to capture individual agents’ strict preferences, such as each agent avoiding obstacles, or avoiding collision with other agents.
-It should be noted that coupling amongst agents’ decisions is captured through constraints in CPTGs.
+The aim of the constraint function $g$ is to capture individual agents’ strict preferences, such as each agent avoiding obstacles or avoiding collision with other agents.
+It should be noted that coupling amongst agents’ decisions is captured through constraints.
 We assume that each agent $i$ aims to minimize an objective function.
@@
 Typically, the reference states for each agent is a path of minimum cost they would follow in the absence of other agents in the environment, e.g., a straight line connecting the start and goal location.
+Such a quadratic cost structure is a standard assumption in various motion planning domains [32].
 Furthermore, we denote the joint reference state of all agents at time $t$ as $\hat{x}_{t}$.
-We use $\mathop{\rm blkdiag}(A_{1},\ldots,A_{n})$ to denote the block-diagonal of the matrices $A_{1},\ldots,A_{n}$.
+We denote by $\mathop{\rm blkdiag}(A_{1},\ldots,A_{n})$ the block-diagonal matrix formed by placing matrices $A_{1},\ldots,A_{n}$ on the diagonal with zeros elsewhere.
 Let $a_{0:\tau}:=\{a_{0},\ldots,a_{\tau}\}$ denote the collection of vectors $a_{t}$’s over the entire horizon $0\leq t\leq\tau$.
-We use $\{{x}_{0:\tau}^{i},{u}_{0:\tau}^{i}\}_{i=1}^{N}$ to denote joint trajectories of all $N$ agents across all times.
+We use $\{{x}_{0:\tau}^{i},{u}_{0:\tau}^{i}\}_{i=1}^{N}$ to denote the joint trajectory of all $N$ the agents at all times.
 Note that using the joint notation, we can equivalently write the joint trajectories of all agents as $\{{x}_{0:\tau},{u}_{0:\tau}\}$.
@@
 Due to the interactive nature of the problem, generally, it is not possible for all agents to minimize their costs simultaneously while satisfying the constraints.
-Therefore, the interaction outcome is best captured by local GNE trajectories of the underlying trajectory game.
+Therefore, the interaction outcome is best captured by local GNE trajectories.
 With dynamics as (1) and constraints as (2), we denote the game by
@@
 Definition 1.
-A set of joint trajectories $\{{x}_{0:\tau}^{i^{\star}},{u}_{0:\tau}^{i^{\star}}\}_{i=1}^{N}$ is a local GNE trajectory for $\mathcal{G}$ if for all agent $i\in\mathcal{N}$, agent i’s trajectory $\{{x}_{0:\tau}^{i^{\star}},{u}_{0:\tau}^{i^{\star}}\}$ is an optimal solution of the following optimization problem withing a local neighborhood $\mathcal{D}\left(\{{x}^{\star}_{0:\tau},{u}^{\star}_{0:\tau}\},\epsilon\right)$ for some $\epsilon>0$:
+A set of joint trajectories $\{{x}_{0:\tau}^{i^{\star}},{u}_{0:\tau}^{i^{\star}}\}_{i=1}^{N}$ is a local GNE trajectory for $\mathcal{G}$ if for every agent $i\in\mathcal{N}$, the agent’s trajectory $\{{x}_{0:\tau}^{i^{\star}},{u}_{0:\tau}^{i^{\star}}\}$ is an optimal solution of the following optimization problem within a local neighborhood $\mathcal{D}\left(\{{x}^{\star}_{0:\tau},{u}^{\star}_{0:\tau}\},\epsilon\right)$ for some $\epsilon>0$:
 $\min_{\{x_{0:\tau}^{i},u_{0:\tau}^{i}\}}$ $\left\lVert x_{\tau}^{i}-\hat{x}_{\tau}^{i}\right\rVert_{Q^{i}_{\tau}}^{2}+\sum_{t=0}^{\tau-1}\left\lVert x_{t}^{i}-\hat{x}_{t}^{i}\right\rVert_{Q^{i}}^{2}+\sum_{t=0}^{\tau}\left\lVert u_{t}^{i}\right\rVert_{R^{i}}^{2}$
@@
 $\tilde{u}_{t}$ $\coloneqq\begin{bmatrix}({u}_{t}^{1^{\star}})^{\top}&({u}_{t}^{2^{\star}})^{\top}&\cdots&({u}_{t}^{{i-1}^{\star}})^{\top}&(u_{t}^{i})^{\top}&({u}_{t}^{{i+1}^{\star}})^{\top}&\cdots&({u}_{t}^{N^{\star}})^{\top}\end{bmatrix}^{\top}$.
-Intuitively, at a local GNE, no agent can reduce their cost function by independently changing their trajectory to any alternative feasible trajectory within a local neighborhood of the joint equilibrium trajectory, $\{{x}_{0:\tau}^{i^{\star}},{u}_{0:\tau}^{i^{\star}}\}_{i=1}^{N}$.
-Note that in CPTGs, the agents’ trajectories are coupled through their joint constraints.
+Intuitively, at a local GNE, no agent can reduce their cost function by independently changing their trajectory to any alternative feasible trajectory within the local neighborhood of the equilibrium trajectories.
 It is important to highlight that solving (1) requires solving a set of $N$ coupled constrained optimal control problems, which is often computationally expensive.
-Therefore, finding all the local GNEs of the game can be time-consuming when using traditional game theoretic methods that aim to solve coupled optimal control problems.
-However, an important property of CPTGs is that a GNE of the original game can be computed by optimizing a single constrained optimal control problem called Potential, as described in [11].
+Therefore, finding all the local GNEs of the game can be time-consuming when using traditional game-theoretic methods that aim to solve coupled optimal control problems.
+However, an important property of our trajectory game is that it is a potential game, i.e., a GNE of the original game can be computed by optimizing a single constrained optimal control problem called Potential, as described in [11].
 This reduces the computational cost of computing the GNEs of the game.
-In the following, we will prove that game $\mathcal{G}$ can be expressed as a constrained trajectory potential game.
-The following theorem shows that any local minimizer of such a potential function is a local generalized Nash equilibrium trajectory.
+In the following, we will prove that game $\mathcal{G}$ can be expressed as a constrained trajectory potential game and that any local minimizer of the potential function is a local GNE.
 Theorem 1.
-A set of joint trajectories $\{{x}_{0:\tau}^{i^{\star}},{u}_{0:\tau}^{i^{\star}}\}_{i=1}^{N}$ is a local GNE trajectory if within a local neighborhood $\mathcal{D}\left(\{{x}^{\star}_{0:\tau},{u}^{\star}_{0:\tau}\},\epsilon\right)$, it is an optimal solution of
+A trajectory $\{{x}_{0:\tau}^{i^{\star}},{u}_{0:\tau}^{i^{\star}}\}_{i=1}^{N}$ is a local GNE if within a local neighborhood $\mathcal{D}\left(\{{x}^{\star}_{0:\tau},{u}^{\star}_{0:\tau}\},\epsilon\right)$, it is an optimal solution of
 $\min_{\{x_{0:\tau},u_{0:\tau}\}}$ $\left\lVert x_{\tau}-\hat{x}_{\tau}\right\rVert_{Q_{\tau}}^{2}+\sum_{t=0}^{\tau-1}\left\lVert x_{t}-\hat{x}_{t}\right\rVert_{Q}^{2}+\sum_{t=0}^{\tau}\left\lVert u_{t}\right\rVert_{R}^{2}$
@@
 $\sum_{i=1}^{N}\left(\left\lVert x_{\tau}^{i}-\hat{x}_{\tau}^{i}\right\rVert_{Q^{i}_{\tau}}^{2}+\sum_{t=0}^{\tau-1}\left\lVert x_{t}^{i}-\hat{x}_{t}^{i}\right\rVert_{Q^{i}}^{2}+\sum_{t=0}^{\tau}\left\lVert u_{t}^{i}\right\rVert_{R^{i}}^{2}\right).$ (6)
-Furthermore, using the joint notations, we can re-write the sums in (6) as (1).
-Therefore, we have proven that $\mathcal{G}$ is a CPTG with potential (1).
+Using the joint notations, we can re-write the sums in (6) as (1).
+Therefore, we have proven that $\mathcal{G}$ is a CPTG with potential function (1).
 Hence, using Theorem 2 from [11], any local solution of (1) will also be a local GNE of $\mathcal{G}$.
 ∎
-Using Theorem 1, any local solutions of (1) will also be local GNE trajectories.
-We would like to acknowledge that finding local solutions of (1) will not necessarily give us all the GNEs of the original as Theorem 1 provides only a sufficient condition for computing local GNEs.
+Using Theorem 1, we want to find all local solutions of (1) that will also be local GNE trajectories as they correspond to different interaction modes and outcomes.
+We would like to acknowledge that finding local solutions of (1) will not necessarily give us all the GNEs of the original, as Theorem 1 provides only a sufficient condition for computing local GNEs.
 Furthermore, it is known that the problem of finding all Nash equilibria is a PPAD-complete problem even in the case of static games [33].
-However, for all practical purposes in multi-agent navigation, finding local solutions of (1) has proven to be enough for us.
-The conventional approach to solving the optimization (1) relies on techniques such as interior point methods [24] or sequential quadratic programming [25].
+However, for practical purposes in multi-agent navigation, finding local solutions of (1) has proven to be enough for us.
+The conventional approach to solving (1) relies on techniques such as interior point methods [24] or sequential quadratic programming [25].
 Naively, the local solutions can be obtained by exhaustive initialization of random trajectories to an optimization solver, such as IPOPT [24], hoping to explore the solution space of local Nash equilibria.
 However, randomly exploring the nonconvex optimization landscape via gradient-based methods is computationally prohibitive and inefficient, as the same local equilibria can be arrived at from multiple random initializations.
-To address this, we utilize a Bayesian inference approach in the following section to efficiently recover all local minima of (1).
+To address this, we utilize a Bayesian inference approach in the following section to efficiently recover all identified minima of (1).
+Figure 2: When using MultiNash-PF on a difficult two agent trajectory game, we first employ an implicit particle filter to obtain coarse estimates of GNEs.
 IV Identifying Local Nash Equilibria via Bayesian Inference
@@
 To this end, we first reformulate problem (1) by considering soft-constraints and introducing auxiliary variables $y_{x,t}$ and $y_{g,t}$ as:
-$\min_{\{v_{t},w_{t},\eta_{t}\}_{t=0}^{\tau}}$ $\left\lVert v_{\tau}\right\rVert_{Q_{\tau}}^{2}+\sum_{t=0}^{\tau-1}\left\lVert v_{t}\right\rVert_{Q}^{2}+\sum_{t=0}^{\tau}\left\lVert w_{t}\right\rVert_{R}^{2}+\left\lVert\eta_{t}\right\rVert^{2}_{Q_{\eta}},$
+$\min_{\{v_{t},w_{t},\eta_{t}\}_{t=0}^{\tau}}$ $\left\lVert v_{\tau}\right\rVert_{Q_{\tau}}^{2}+\sum_{t=0}^{\tau-1}\left\lVert v_{t}\right\rVert_{Q}^{2}+\sum_{t=0}^{\tau}\left(\left\lVert w_{t}\right\rVert_{R}^{2}+\left\lVert\eta_{t}\right\rVert^{2}_{Q_{\eta}}\right),$
 $\mathrm{s.t}\quad$ $\begin{bmatrix}x_{t+1}\\ u_{t+1}\end{bmatrix}=\begin{bmatrix}f(x_{t},u_{t})\\ 0_{m}\end{bmatrix}+\begin{bmatrix}0_{n}\\ w_{t}\end{bmatrix}$,
@@
 $\begin{bmatrix}x_{0}\\ u_{0}\end{bmatrix}=\begin{bmatrix}\hat{x}_{0}\\ w_{t-1}\end{bmatrix},\quad 0\leq t\leq\tau,$ (7)
-where $v_{t}$, $w_{t}$, and $\eta_{t}$ are bounded disturbances, and $Q_{\eta}\in\mathbb{R}^{n_{c}\times n_{c}}$ assigns the importance of constraint satisfaction.
-Further, $\psi$ is a barrier function used to quantify the degree of constraint violation by defining
+where $v_{t}$, $w_{t}$, and $\eta_{t}$ are bounded disturbances.
+The weight matrix $Q_{\eta}\in\mathbb{R}^{n_{c}\times n_{c}}$ quantifies the slackness in satisfying constraints.
+Further, the degree of constraint violation is captured by a barrier function $\psi(\cdot,\cdot)$, which is defined as
 $\psi\left(g(x_{t},u_{t})\right)=\frac{1}{\alpha}\ln\left(1+\exp\left(g\left(x_{t},u_{t}\right)\right)\right),$ (8)
-where $\alpha$ controls the strictness enforcing constraints.
-We note that the problem in (IV) resembles a moving horizon estimation problem [32].
+where $\alpha$ is a parameter that tunes the strictness of constraint enforcement.
 A direct attempt at solving (IV) using traditional gradient-based methods poses a similar computational burden in identifying all of the local equilibria compared to problem (1).
-However, the above connection between (1) and the moving horizon estimation problem (IV) highlights the potential to treat the problem from a Bayesian inference perspective.
+However, the above connection between (1) and the moving horizon estimation problem (IV) highlights the possibility to view the problem from a Bayesian inference perspective.
 To this end, we construct the following virtual state-space model from (IV) as
@@
 The value of the virtual measurements must be set in a way that steers the inference problem to have the same optima as in problem (1).
-We achieve this by noting that $y_{x,t}$ is a measurement value used to correct the state estimate of $x_{t}$ towards the reference $\hat{x}_{t}$.
-While $y_{g,t}$ is to satisfy constraints through (8).
-Observing that the barrier function in (8) outputs zero when the constraints are satisfied, we set the virtual measurement for the barrier function equal to $0_{c}$.
+We achieve this by noting that $y_{x,t}$ is a measurement value used to correct the state estimate of $x_{t}$ towards the reference $\hat{x}_{t}$ while $y_{g,t}$ is to satisfy constraints through (8).
+Observing that the barrier function in (8) outputs zero when the constraints are satisfied, we set the virtual measurement for the barrier function to be equal to $0_{c}$.
 Hence, the virtual measurement that will drive $\bar{x}_{t}$ towards optimality takes the value
@@
 The Bayesian inference problem pertaining to (9) is one that characterize the conditional distribution $p(\bar{x}_{0:\tau}\,|\,\bar{y}_{0:\tau})$.
-In the following, lemma (1) elucidates that equivalence holds between a posteriori estimate of $p(\bar{x}_{0:\tau}\,|\,\bar{y}_{0:\tau})$ and the problem in (1).
+In the following lemma, we show that an equivalence holds between a posteriori estimate of $p(\bar{x}_{0:\tau}\,|\,\bar{y}_{0:\tau})$ and (1).
 Lemma 1.
@@
 $\bar{x}_{0:\tau}^{\star}=\arg\max_{\bar{x}_{0:\tau}}\>\log p(\bar{x}_{0:\tau}\,|\,\bar{y}_{0\tau}).$
-We refer the interested reader to [16, Thm.
+We refer the interested reader to [14, Thm.
 1] for the proof.
-The above lemma implies that the local GNEs in (1) translate to multiple local maximums of $p(\bar{x}_{0:\tau}\,|\,\bar{y}_{0:\tau})$.
+The above lemma implies that the local GNEs in (1) translate to multiple local maxima of $p(\bar{x}_{0:\tau}\,|\,\bar{y}_{0:\tau})$.
 Hence, the Bayesian inference problem for (9) involves characterizing the multi-modal PDF $p(\bar{x}_{0:\tau}\,|\,\bar{y}_{0:\tau})$.
@@
 where $\delta(\cdot)$ is the Dirac delta function and $w_{\tau}^{i}$ is the weight assigned to a sample trajectory $\bar{x}_{0:\tau}^{i}$ for $j=1,\ldots,J$.
-To achieve the approximation in (10), we employ the implicit particle filter  [31] due to the twofold benefit it holds to implement the MultiNash-PF framework.
-First, as suggested in lemma 1, the local maximums (i.e., high-probability regions) of $p(\bar{x}_{0:\tau}\,|\,\bar{y}_{0:\tau})$ correspond to the local GNEs.
+It is important to note that, in our setting, each particle represents a complete trajectory.
+Figure 3: MultiNash-PF identifies all the local GNE trajectories for a challenging two-agent trajectory game where two agents swapped positions while avoiding an obstacle.
+In this scenario, the obstacle radius is greater than the collision avoidance radius of the two agents, leading to the discovery of 6 distinct Nash equilibria.
+MultiNash-PF effectively obtains coarse estimates of local GNEs through implicit particle filtering, which are then refined using IPOPT to obtain local GNEs.
+To achieve the approximation in (10), we employ the implicit particle filter in [31] due to the twofold benefit it holds.
+First, as suggested in Lemma 1, the local maxima (i.e., high-probability regions) of $p(\bar{x}_{0:\tau}\,|\,\bar{y}_{0:\tau})$ correspond to the local GNEs.
 Hence, focusing our sampling efforts towards these high-probability regions can reduce the number of required samples for the approximation in (10).
-This approach aligns precisely with the principle of implicit importance sampling, whose objective is to build a mapping hat draws particles from the high-probability regions of the filtering distribution in (10).
-Second, we want to avoid the traditionally needed resampling step in particle filtering methods to ensure that the inferred trajectories are smooth.
-Fortunately, the implicit particle filter does not suffer from the problem of particle degeneracy, and filtering can be achieved by disregarding resampling.
-We leverage the Markovian property of the virtual system (9) to recursively approximate (10). Using Bayes’ rule, we have
+This approach aligns with the principle of implicit importance sampling, whose objective is to build a mapping that draws particles from the high-probability regions of the filtering distribution in (10).
+The implicit particle filter mitigates the particle degeneracy problem that pertains to particle filters by focusing the samples on the high-probability regions, which results in samples with large enough weights that are less prone to degeneration.
+Therefore, the implicit particle filter can often be deployed without requiring a resampling step [31], thereby resulting in smoother approximated trajectories, although occasional resampling may still be required.
+We leverage the Markovian property of the virtual system (9) to recursively approximate (10) using Bayes’ rule:
 $p(\bar{x}_{0:t}\,|\,\bar{y}_{0:t})=p(\bar{y}_{t}\,|\,\bar{x}_{t})p(\bar{x}_{t}\,|\,\bar{x}_{t-1})p(\bar{x}_{0:t-1}\,|\,\bar{y}_{0:t-1}),$ (11)
@@
 This indicates that given the particle $\bar{x}_{t-1}^{j}$ from the prior distribution $p(\bar{x}_{0:t-1}\,|\,\bar{y}_{0:t-1})$, we can obtain $\bar{x}_{t}^{j}$.
-As shown in [31], we use implicit importamapping sampling and identify an implicit mapping to sample from high probability regions of (11).
-The implicit map is constructed by considering sampling from a reference random vector $\xi_{t}$ with density $p(\gamma_{t})$ and map it to high probability regions of $p(x_{0:\tau}\,|\,y_{0:\tau})$, such that a sample $\gamma^{i}_{t}$ is mapped to $x_{t}^{i}$.
+As shown in [31], we use implicit importance sampling and identify an implicit mapping to sample from high probability regions of (11).
+The implicit map is constructed by considering sampling from a reference random vector $\xi_{t}$ with density $p(\gamma_{t})$ and mapping it to high probability regions of $p(x_{0:\tau}\,|\,y_{0:\tau})$, such that a sample $\gamma^{i}_{t}$ is mapped to $x_{t}^{i}$.
 The mapping connects the highly probable regions of $p(\xi)$ to $p(x_{0:\tau}\,|\,y_{0:\tau})$ by letting
@@
 Finding an explicit solution to (12) is computationally intractable due to the involved nonconvex optimization.
-To bypass this computational challenge, a local Gaussian approximation around $\bar{x}_{t-1}^{j}$ is employed such that
+To bypass this computational challenge, we employ a local Gaussian approximation around $\bar{x}_{t-1}^{j}$ such that
 $p(\bar{x}_{t}^{j}|\bar{y}_{0:t})\sim\mathcal{N}\left(\hat{m}_{t}^{j},\hat{\Sigma}_{t}^{x,j}\right),$
@@
 where $\gamma_{t}^{j}$ is a sample from $p(\gamma_{t})\sim\mathcal{N}(0,I_{n})$ and $I_{n}$ denotes the identity matrix of size $n$.
-This process is repeated until $t=\tau$ for each particle to obtain the set of trajectories $\{\bar{x}_{0:\tau}\}_{j=1}^{J}$, which will include the coarse estimates of local GNEs.
-This completes the unscented implicit particle filter, identifying the coarse estimates of multiple local Nash equilibria in (1).
-In the last step, the local equilibria are extracted to be from $\{x_{0:\tau}\}_{j=1}^{J}$.
+We repeat this process until $t=\tau$ for each particle to obtain the set of trajectories $\{\bar{x}_{0:\tau}\}_{j=1}^{J}$, which will include the coarse estimates of local GNEs.
+This completes the unscented implicit particle filter, which identifies the coarse estimates of multiple local Nash equilibria in (1).
+In the last step, we extract local equilibria from $\{x_{0:\tau}\}_{j=1}^{J}$.
 As viewed in Fig.
-4, the particle set contains multiple trajectories for each GNE.
-Therefore, to isolate the GNE trajectories, a clustering method is used to aggregate the particles close to each other in one cluster.
-One can leverage methods such as hierarchical clustering [35] or DBSCAN [36].
+2, the particle set contains multiple trajectories for each GNE.
+Therefore, to isolate the GNE trajectories, we use a clustering method to aggregate the particles close to each other in one cluster.
+We leverage methods such as hierarchical clustering [36] or DBSCAN [37] for clustering purposes.
 It should be noted that clustering methods require a metric to compute distances between the data points.
-The Fréchet distance [37] is utilized as our metric to compute distances between the trajectories.
-The clustering of trajectories will result in one cluster for each GNE by taking the average of trajectories over each cluster and feed them as a warm start to the IPOPT solver to obtain the local GNEs.
+We use Fréchet distance [38] as our metric to compute distances between the trajectories.
+The clustering of trajectories will result in one cluster for each GNE.
+For each cluster, we compute the average trajectory which serves as an initial guess for the IPOPT solver.
+The solver then refines this initial guess by minimizing (1), ultimately converging to a precise local GNE.
 The algorithmic steps of the proposed MultiNash-PF are provided in Algorithm 1.
@@
 3:Perform unscented implicit particle filtering as follows:
-4:Initialize the particles: $\bar{x}_{0}^{j}$, $\hat{\Sigma}^{x,j}_{0}$, and $w_{0}^{j}$ for $j=1,\ldots,J$
-5:Employ Algorithm-1 from [31] combined with UKF to obtain coarse solutions, $\{x_{0:\tau}\}_{j=1}^{J}$.
-6:Extract trajectories through clustering from $\{x_{0:\tau}\}_{j=1}^{J}$ and warm start IPOPT solver to obtain the local GNEs $\{{x}_{0:\tau}^{i^{\star}},{u}_{0:\tau}^{i^{\star}}\}_{i=1}^{N}$ of $\mathcal{G}$.
+[…]
+6:  Employ Algorithm-1 from [31] combined with UKF
+7:  to obtain coarse solutions, $\{x_{0:\tau}\}_{j=1}^{J}$.
+8:Extract trajectories through clustering from $\{x_{0:\tau}\}_{j=1}^{J}$ and warm start IPOPT solver to obtain the local GNEs $\{{x}_{0:\tau}^{i^{\star}},{u}_{0:\tau}^{i^{\star}}\}_{i=1}^{N}$ of $\mathcal{G}$.
 V Numerical Simulations
@@
 We consider the following discrete-time unicycle dynamics to model each agent $i\in\{1,2\}$:
-$p^{i}_{t+1}$ $=p^{i}_{t}+\Delta t\cdot v^{i}_{t}\cos{\theta^{i}_{t}},\;q^{i}_{t+1}=q^{i}_{t}+\Delta t\cdot v^{i}_{t}\sin{\theta^{i}_{t}}$
-$\theta^{i}_{t+1}$ $=\theta^{i}_{t}+\Delta t\cdot\omega_{i},\;v^{i}_{t+1}=v^{i}_{t}+\Delta v^{i}_{t},\;\omega^{i}_{t+1}=\omega^{i}_{t}+\Delta\omega^{i}_{t},$
-where $\Delta t$ is the time-step size, $p^{i}_{t}$ and $q^{i}_{t}$ are $x$ and $y$ coordinates of the positions, $\theta^{i}_{t}$ is the heading angle from positive x-axis, $v^{i}_{t}$ is the linear velocity, and $\omega^{i}_{t}$ is the angular velocity of the agent $i$ at time $t$.
-Further, $\Delta v^{i}_{t}$ is the change in linear velocity, and $\Delta\omega^{i}_{t}$ is the change in angular velocity of the agent $i$ at time $t$, respectively.
+$p^{i}_{t+1}$ $=p^{i}_{t}+\Delta t\cdot\nu^{i}_{t}\cos{\theta^{i}_{t}},\;q^{i}_{t+1}=q^{i}_{t}+\Delta t\cdot\nu^{i}_{t}\sin{\theta^{i}_{t}}$
+$\theta^{i}_{t+1}$ $=\theta^{i}_{t}+\Delta t\cdot\omega_{i},\;\nu^{i}_{t+1}=\nu^{i}_{t}+\Delta\nu^{i}_{t},\;\omega^{i}_{t+1}=\omega^{i}_{t}+\Delta\omega^{i}_{t},$
+where $\Delta t$ is the time-step size, $p^{i}_{t}$ and $q^{i}_{t}$ are $x$ and $y$ coordinates of the positions, $\theta^{i}_{t}$ is the heading angle from positive x-axis, $\nu^{i}_{t}$ is the linear velocity, and $\omega^{i}_{t}$ is the angular velocity of the agent $i$ at time $t$.
+Further, $\Delta\nu^{i}_{t}$ is the change in linear velocity, and $\Delta\omega^{i}_{t}$ is the change in angular velocity of the agent $i$ at time $t$, respectively.
 For each agent $i$, we denote the state vector $x^{i}_{t}$ and the action vector $u^{i}_{t}$ as follows
-$x^{i}_{t}=\begin{bmatrix}p^{i}_{t}&q^{i}_{t}&\theta^{i}_{t}&v^{i}_{t}&\omega^{i}_{t}\end{bmatrix}^{\top},\enskip u^{i}_{t}=\begin{bmatrix}\Delta v^{i}_{t}&\Delta\omega^{i}_{t}\end{bmatrix}^{\top}.$
+$x^{i}_{t}=\begin{bmatrix}p^{i}_{t}&q^{i}_{t}&\theta^{i}_{t}&\nu^{i}_{t}&\omega^{i}_{t}\end{bmatrix}^{\top},\enskip u^{i}_{t}=\begin{bmatrix}\Delta\nu^{i}_{t}&\Delta\omega^{i}_{t}\end{bmatrix}^{\top}$.
 We formulate this scenario as a two-player game according to cost functions as described in (3).
 The reference trajectory for both agents is given as a straight line on the highway, starting from their initial position to their respective goal location.
-With $\hat{Q}=diag(50,10,5,5,2)$, the reference tracking and control effort cost matrices are set as
+With $\hat{Q}=\mathop{\rm blkdiag}(50,10,5,5,2)$, the reference tracking and control effort cost matrices are set as
 $Q^{i}=0.6\hat{Q},\;Q^{i}_{\tau}=100\hat{Q},\;R^{i}=diag(8,4).$ (13)
@@
 $g_{c}(x_{t},u_{t})\coloneq-d(x^{1}_{t},x^{2}_{t})+r_{\text{col}}\leq 0,$ (14)
-where $d(x^{1}_{t},x^{2}_{t})\coloneq\sqrt{(p^{1}_{t}-p^{2}_{t})^{2}+(q^{1}_{t}-q^{2}_{t})^{2}}$ is the Euclidean distance between the two agents and $r_{\text{col}}\coloneq 3m$ is the collision avoidance radius between the two agents.
+where $d(x^{1}_{t},x^{2}_{t})$ is the Euclidean distance between the two agents and $r_{\text{col}}\coloneq 3m$ is the collision avoidance radius between the two agents.
 With this setup, we employ MultiNash-PF from Algorithm 1 to efficiently capture the multi-modality in the solution space and then extract the solutions using hierarchical clustering [36] to provide a warm start trajectory to IPOPT [24] solver to recover different local GNEs.
@@
 As we can see, the presence of two agents and collision avoidance constraints induces two different possible GNEs.
-Either both the agents will yield to their right, or they will yield to their left to avoid collision.
+Either both agents will yield to their right, or they will yield to their left to avoid collision.
 As shown in Fig 4, our method can recover both equilibria.
-Furthermore, we consider a more complicated scenario in which two agents aim to exchange their positions while avoiding an obstacle of radius $r_{\text{obs}}=4m$ located at the midpoint of the two agents.
-It should be noted that the radius of the obstacle is larger than the collision avoidance radius of the two agents.
-We consider the same unicycle dynamics as before, and the reference trajectory cost is also the same as in (13).
-Furthermore, the collision avoidance radius between the two agents is also the same as before.
-In addition to (14), the constraints for this game will include both agents avoiding the obstacle $g^{i}_{o}(x_{t},u_{t}):=\begin{bmatrix}-d(x^{i}_{t},x_{\text{obs}})+r_{\text{obs}}\end{bmatrix}$ for $i=1,2$ where $x_{\text{obs}}$ is the location of the obstacle.
-We also assume that we have constraints on the control bounds $g_{b}(x_{t},u_{t})=\begin{bmatrix}-v^{1}_{t},&-v^{2}_{t},&|u^{1}_{t}|-u^{1}_{b},&|u^{2}_{t}|-u^{2}_{b}\end{bmatrix}$, where $u^{1}_{b}$ and $u^{2}_{b}$ are controlled action bounds for agents 1 and 2, respectively.
-We write the combined constraints as $g(x_{t},u_{t}):=\begin{bmatrix}g_{c}&g^{1}_{o}&g^{2}_{o}&g_{b}\end{bmatrix}^{\top}\leq 0.$ We choose the values of control action bounds to be $u^{1}_{b}=u^{2}_{b}=[0.15,0.75]^{\top}$.
+Figure 4: The Nash equilibrium trajectories found by MultiNash-PF when two unicycle agents change their positions.
+First, we use the implicit particle filter to discover the coarse estimates of two different equilibria, and then utilize them as a warm start for IPOPT to obtain the two Nash equilibrium trajectories.
+Next, we consider a more complicated scenario in which two agents aim to exchange their positions while avoiding an obstacle of radius $r_{\text{obs}}=4m$ located at the midpoint of the two agents.
+Note that the radius of the obstacle is larger than the collision avoidance radius of the two agents.
+We consider the same setup as before.
+In this example, in addition to (14), the constraints for this game will include both agents avoiding the obstacle $g^{i}_{o}(x_{t},u_{t}):=\begin{bmatrix}-d(x^{i}_{t},x_{\text{obs}})+r_{\text{obs}}\end{bmatrix}$ for $i=1,2$ where $x_{\text{obs}}$ is the location of the obstacle.
+We also assume that we have constraints on the control bounds $g_{b}(x_{t},u_{t})=\begin{bmatrix}-\nu^{1}_{t},&-\nu^{2}_{t},&|u^{1}_{t}|-u^{1}_{b},&|u^{2}_{t}|-u^{2}_{b}\end{bmatrix}$, where $u^{1}_{b}$ and $u^{2}_{b}$ are controlled action bounds for agents 1 and 2, respectively.
+We write the combined constraints as $g(x_{t},u_{t}):=\begin{bmatrix}g_{c}&g^{1}_{o}&g^{2}_{o}&g_{b}\end{bmatrix}^{\top}\leq 0.$ We choose the values of control input bounds to be $u^{1}_{b}=u^{2}_{b}=[0.15,0.75]^{\top}$.
 This setup induces multitudes of local equilibria.
@@
 It is challenging to recover all six equilibria without utilizing the capabilities of the implicit particle filter.
+We fixed $J=50$ particles and obtained the solution trajectories using implicit particle filtering from Algorithm 1 described in Section IV.
 Upon performing simulation experiments with our method, we observe that MultiNash-PF can recover all six Nash equilibria as shown in Fig.
@@
 3.
-(a) Histogram of the solve time comparison of MultiNash-PF and the baseline.
-(b) Histogram of the number of the IPOPT runs required to obtain all six equilibria.
+Please note that, while trajectories may overlap spatially, collisions are avoided because agents traverse those regions at different times, as enforced by the constraints in our optimization.
+(a) Histogram of the number of IPOPT runs required to obtain all six equilibria.
+(b) Histogram of the solve time comparison of MultiNash-PF and the baseline.
 Figure 5: Comparison of MultiNash-PF with the baseline for the experiment shown in Fig. 3.
-(a) The average solve time for MultiNash-PF is $3.36\pm 0.59s$ while the baseline’s average solve time is $6.71\pm 2.83s$.
-MultiNash-PF reduces the solve time by about $50\%$ compared to the baseline while giving a much lower variance on the solve time.
-(b) For the baseline, the average number of IPOPT runs required to obtain all six different equilibria is $18.64\pm 7.57$, while for MultiNash-PF, this is always $6$.
+(a) For the baseline, the average number of IPOPT runs required to obtain all six different equilibria is $18.64\pm 7.57$, while for MultiNash-PF, this is always $6$.
 MultiNash-PF requires 3 times fewer IPOPT runs.
+(b) MultiNash-PF reduces the solve time by about $50\%$ compared to the baseline while giving a much lower variance on the solve time.
 V-1 Comparison with Baseline
 In order to compare the effectiveness of our method, we consider a baseline method.
-In the baseline method, we directly initialize the IPOPT solver with random perturbations of reference trajectories and compute the resulting optimal solutions.
-For our method, we fixed $J=50$ particles and obtained the solution trajectories using implicit particle filtering from Algorithm 1 described in Section IV.
-After that, we extract all the different equilibria and provide them as a warm starting point for IPOPT to obtain the final solutions.
-The compute time of our method is a combination of time used in implicit particle filtering and time used by IPOPT to obtain all the different equilibria given warm starting from filtering.
-We run the IPOPT solver until we recover all 6 different equilibria.
+For the baseline, we directly initialize the IPOPT solver with random perturbations of reference trajectories and compute the resulting optimal solutions.
+The computation time of our method is a combination of the time used in implicit particle filtering and the time used by IPOPT to obtain all the different equilibria given warm starting from filtering.
+For the baseline, we randomly initialize the IPOPT solver until we recover all 6 different equilibria.
 We repeat this experiment for 100 different Monte Carlo runs.
-We observe that it requires, on average, $18.64\pm 7.57$ number of IPOPT runs to obtain all six different equilibria using the baseline approach of random warm-starting.
+We observe that it requires, on average, $18.64\pm 7.57$ random initializations of IPOPT to obtain all six different equilibria using the baseline approach.
 However, with our method, the number of IPOPT runs required is always six because we always obtain all six modes of equilibria from implicit particle filtering.
 A histogram of the number of IPOPT runs required for obtaining all six equilibria is plotted in Fig.
-5(b).
+5(a).
 Furthermore, the solve time to recover all 6 different Nash equilibria for the baseline is $6.71\pm 2.83s$, while for MultiNash-PF, the solve time is $3.36\pm 0.59s$.
@@
 A histogram of the solve times of the baseline compared to our method is plotted in Fig.
-5(a).
+5(b).
 As explained earlier, our method’s solve time is a combination of the time used in implicit particle filtering and the time used by IPOPT.
 We observe that the time required for implicit particle filtering over 100 Monte Carlo runs is $0.37\pm 0.072s$ while the time used by IPOPT to obtain the final solutions is $2.99\pm 0.57s$.
-Therefore, in our MultiNash-PF, the time used by implicit particle filtering is $11\%$ of the total solve time, which shows that filtering is very efficient in exploring the solution space of local Nash equilibria and contributes really less to the overall computation time of MultiNash-PF.
+Therefore, in our MultiNash-PF, the time used by implicit particle filtering is $11\%$ of the total solve time, which shows that filtering is very efficient in exploring the solution space of local Nash equilibria.
 VI Hardware Experiment
-Finally, we conduct a hardware experiment where a human and a robot exchange their position as shown in Fig.
+Finally, we conduct a hardware experiment where a human and a robot exchange their positions as shown in Fig.
 1.
-We perform these experiments using a TurtleBot 4 which is a ground robot with unicycle dynamics.
+We perform these experiments using a TurtleBot 4, which is a ground robot with unicycle dynamics.
 We assume that the human also follows unicycle dynamics.
-The collision avoidance constraint between the TurtleBot and human is assumed to be the same as of (14).
+The collision avoidance constraint between the TurtleBot and the human is assumed to be the same as that of (14).
 We use the Vicon motion capture system to measure the current position of both the robot and the human.
 Similar as shown in Fig.
-2, two GNEs may exist which will correspond to two different modes of interactions.
-Namely, either the human and robot yield to their right or they yield to their left.
+4, two GNEs may exist, which will correspond to two different modes of interactions.
+Namely, yield-right equilibrium when the human and robot yield to their right and yield-left equilibrium when they yield to their left.
 For the robot to safely navigate around the human without collision, the robot needs to be aware of both the GNEs and needs to be able to determine which GNE the human chooses.
-Using MultiNashPF, the robot first uses an implicit particle filter to obtain the coarse estimates of the two different equilibria and then utilizes them as a warm start for IPOPT to obtain two modes of interaction.
+Using MultiNash-PF, the robot first uses an implicit particle filter to precompute the coarse estimates of the two different equilibria and then utilizes them as a warm start for IPOPT to obtain two modes of interaction.
 We assume that the interaction happens over $T=10s$.
-Then similar to [11], the robot uses a potential function based game theoretic planner in a model-predictive fashion with a planning horizon of 5s and $\Delta t=0.1s$ to plan its motion.
-The robot faces two choices of GNEs to follow recovered from the MultiNashPF.
+Then, similar to [11], the robot uses a potential function based game-theoretic planner in a model-predictive fashion with a planning horizon of 5s and $\Delta t=0.1s$ to plan its motion.
+The robot faces two precomputed choices of GNEs to follow recovered from the MultiNash-PF.
 At any given time $t$ of planning, the robot computes the Fréchet distance [38] between the trajectory exhibited by the human so far and the human’s trajectory in both the computed GNEs.
-Let the distances of human trajectory from both the GNEs be denoted by $d_{1}$ and $d_{2}$, respectively.
-The robot keeps an uncertainty on the GNE that the human chooses until $|d_{1}-d_{2}|>d^{*}$ where $d^{*}$ is the specified threshold.
+Let the distances of the human trajectory from both the GNEs be denoted by $d_{1}$ and $d_{2}$, respectively.
+The robot keeps an uncertainty on the GNE that the human chooses until $|d_{1}-d_{2}|>d^{*}$ where $d^{*}$ is a pre-specified threshold.
 For example, in Fig.
-1 at $t=2$, values of $d_{1}$ and $d_{2}$ are almost identical which indicates that human has not chosen their mode of interaction.
+1, at $t=2$, the values of $d_{1}$ and $d_{2}$ are almost identical, which indicates that the human has not chosen their mode of interaction.
 Whenever $|d_{1}-d_{2}|>d^{*}$ occurs, the robot becomes certain of the mode of interaction followed by the human.
-Then the robot follows GNE corresponding to $\arg\min_{i\in\{1,2\}}\{d_{i}\}$.
+Then, the robot follows the GNE corresponding to $\arg\min_{i\in\{1,2\}}\{d_{i}\}$.
 For example, in Fig.
-1, at $t=4$, the Fréchet distance of the human from its first GNE trajectory is very high while for the second GNE, the Fréchet distance is low indicating that the human has chosen the second GNE which corresponds to both the agents yielding to their right.
+1, at $t=4$, the Fréchet distance of the human from its yield-left equilibrium trajectory is very high, while for the yield-left equilibrium, the Fréchet distance is low, indicating that the human has chosen the yield-right equilibrium.
 Using our method, the robot autonomously identifies this mode, moves accordingly, and reaches its goal location while successfully avoiding collision with the human.
-We ran 10 different trials and each time we asked the human to either yield to their right or left randomly.
+We ran 10 different trials, and each time, we asked the human to either yield to their right or left randomly.
 Our algorithm was able to successfully compute both modes and identify the mode that the human was following.
@@
 VII Conclusion
-In this work, we introduced MultiNash-PF, a novel algorithm for efficiently computing multiple local GNEs in CPTGs.
-By leveraging potential game theory and implicit particle filtering, MultiNash-PF efficiently identifies coarse estimates of local minimizers, which are then refined using optimization solvers to obtain distinct interaction modes.
-Our numerical simulations demonstrate that MultiNash-PF significantly reduces computation time—by up to 50% compared to baseline methods—while effectively capturing the multimodal nature of multi-agent interactions.
+In this work, we introduced MultiNash-PF, a novel algorithm for efficiently computing multiple interaction modes.
+By leveraging potential game theory and implicit particle filtering, MultiNash-PF efficiently identifies coarse estimates of interaction modes, which are then refined using optimization solvers to obtain distinct interaction modes.
+Our numerical simulations demonstrate that MultiNash-PF significantly reduces computation time by up to 50% compared to baseline methods, while effectively capturing the multimodal nature of multi-agent interactions.
 Furthermore, our real-world human-robot interaction experiments highlight the algorithm’s ability to reason about multiple interaction modes and resolve conflicts in real-time.
-These results underscore the potential of MultiNash-PF to enhance interactive motion planning, enabling more adaptive and socially aware autonomous systems.